### v2

# ModelMissSol-15-11-23-supp

Abstract
We provide a simple and general solution for accurate uncertainty quantification of Bayesian inference in misspecified or approximate models, and for generalized posteriors more generally. While existing solutions are based on explicit Gaussian posterior approximations, or post-processing procedures, we demonstrate that correct uncertainty quantification can be achieved by substituting the usual posterior with an intuitively appealing alternative posterior that conveys the same information. This solution applies to both likelihood-based and loss-based posteriors, and we formally demonstrate the reliable uncertainty quantification of this approach. The new approach is demonstrated through a range of examples, including linear models, and doubly intractable models.
Keywords: Bayesian inference, Generalized Bayesian inference, Gibbs posteriors, Model misspecification, Calibration.
1 Introduction
Bayesian methods are lauded for their ability to tackle complicated models, deftly handle latent variables, and for providing a holistic treatment for the uncertainty of model unknowns. While it is well-know that Bayesian methods deliver reliable inferences in well-specified models, when the model used to define the posterior is misspecified, the shortcomings of the Bayes posteriors are well-known and much recent literature has been devoted to correcting these issues. One popular strategy for delivering reliable Bayesian inference in misspecified models is to instead produce posteriors belies using a “Gibbs” posteriors based on losses that connect a parameter of interest to observed data (see, e.g., Chernozhukov and Hong, , 2003, Zhang, , 2006, and Bissiri et al., , 2016 for specific architecture); see, e.g., Syring and Martin, (2020), Matsubara et al., (2022), Jewson and Rossell, (2022), and Loaiza-Maya et al., (2021) for specific examples.
Unfortunately, as discussed by several authors, see, e.g., Syring and Martin, (2019) and Miller, (2021), Gibbs posteriors are not calibrated. Following Rubin, (1984), a posterior is said to be calibrated if their “posterior probability statements have the asserted coverage in repeated experience.” Calibration of Gibbs posteriors is crucial if one wishes to make scientific probability statements that can be reliably refuted through empirical analysis (Rubin, , 1984); see, also, Dawid, (1982), and Little, (2006) for additional discussion on the importance of calibrated Bayesian inference.
Several approaches have been suggested to solve this issue (see, e.g., Müller, , 2013, Holmes and Walker, , 2017, Lyddon et al., , 2019, Syring and Martin, , 2019, and Matsubara et al., , 2022, and see, Wu and Martin, , 2020 for a review). However, all the proposed approaches of which we are aware amount to applying extrinsic corrections to some Bayesian belief update.
Herein, our focus is to propose a novel Gibbs posterior that delivers calibrated inferences in settings where generalized, or Gibbs, psoteriors are most commonly applied: settings where correct model specification is in doubt, but where a convenient loss defined by a low-to-moderate dimensional unknown parameter is available for us to base our inferences on. In these settings, we show that our Gibbs posterior delivers calibrated inferences without the need of ad-hoc corrections, or hyper-parameter tuning. Further, this approach is applicable to any class of loss function that are sufficiently smooth in expectation. This new Gibbs posterior allows the “statistician to be Bayesian in principle and calibrated to the real world in practice” (Rubin, , 1984).
The remainder of the paper proceeds as follows. Section 2 discusses the general issue of model misspecification in likelihood-based Bayesian inference, and Section 2.2 demonstrates how a particular Gibbs posterior approach overcomes the known issues with Bayesian inference in this setting. This section also includes two preliminary examples demonstrating the behavior of the proposed method. Section 3 examines the empirical performance of the proposed approach in exact and generalized Bayesian inference paradigms, including two examples of doubly intractable models, and a situation where a computationally convenient approximating model is used instead of the actual model. In each example, our proposed approach delivers reliable and correctly calibrated Bayesian inferences. Section 4 proves that the proposed approach correctly quantifies uncertainty. Section 5 compares the new approach with existing approaches that attempt to produce reliable Bayesian inferences in these settings and Section 6 concludes. Online supplementary material contains proofs of all stated results.
2 Bayesian Inference in Misspecified Models
2.1 Setup and Known Issues
The observed data is $y:=(y_{1},\dots,y_{n})^{\top}$, where $y_{i}\in\mathcal{Y}\subseteq\mathbb{R}^{d_{y}}$ ($i=1,\dots,n$), is generated from some true unknown probability distribution $P^{(n)}_{0}$. Generally, in Bayesian infernece one approximates the unknown $P^{(n)}_{0}$ using a class of models $\{P^{(n)}_{\theta}:\theta\in\Theta\}$ depending on the unknown parameter $\theta\in\Theta\subseteq\mathbb{R}^{d_{\theta}}$, for which we have prior beliefs $\pi(\theta)$, and where the number of parameters $d_{\theta}$ is fixed with $n$.
However, when the model $P^{(n)}_{\theta}$ is misspecified, it is well-known that uncertainty quantification using the exact Bayesian posterior is not reliable (Kleijn and van der Vaart, , 2012), and in such cases the very parameters we are conducting inference on may be “meaningless - except perhaps as descriptive statistics” (Freedman, , 2006). When $P^{(n)}_{\theta}$ is possibly misspecified, it may instead make sense to leave behind the likelihood and consider classes of loss functions that are specific for the inferential task at hand. In such cases, so long as one is willing to take a more general approach to updating prior beliefs, we can follow the ideas of Zhang, (2006), Bissiri et al., (2016), and Knoblauch et al., (2022) to produce a “generalized posterior” by solving an infinite-dimensional optimization problem. Let $\mathcal{D}_{n}:\Theta\rightarrow\mathbb{R}$ denote a loss function associated with the sample $y$, and that depends on $\theta\in\Theta$. We can then produce a “posterior”, or belief update, based on the loss $\mathcal{D}_{n}(\theta)$ by solving the optimization problem
$\pi(\theta\mid\mathcal{D}_{n}):=\operatornamewithlimits{argmin\,}_{\rho\in\mathcal{P}(\Theta)}\left\{\omega\int_{\Theta}\mathcal{D}_{n}(\theta)\rho(\theta)\mathrm{d}\theta+\text{KL}(\rho\|\pi)\right\}.$ (1)
In equation (1), the scaling parameter $\omega\geq 0$, often called the learning rate, scales the loss $\mathcal{D}_{n}(\theta)$ relative to the prior $\pi(\theta)$, and determines the relative weight each term is given in the belief update, with small $\omega$ delivering beliefs that resemble the prior; [^1] The weight $\omega$ serves as a tuning parameter that balances the importance of minimizing $\int_{\Theta}\mathcal{D}_{n}(\theta)\rho(\theta)\mathrm{d}\theta$ relative to matching the prior in KL-divergence. $\text{KL}(\rho\|\pi)$ denotes the Kullback-Liebler (KL) divergence between the measure we are optimizing over, $\rho\in\mathcal{P}(\Theta)$, and the prior $\pi\in\mathcal{P}(\Theta)$, and $\mathcal{P}(\Theta)$ denotes a set of probability measures over $\Theta$. When the loss is integrable with respect to the prior $\pi(\theta)$, the unique solution to (1) takes the recognizable form:
$\pi(\theta\mid\mathcal{D}_{n}):=\frac{\pi(\theta)\exp\{-\omega\cdot\mathcal{D}_{n}(\theta)\}}{\int_{\Theta}\pi(\theta)\exp\{-\omega\cdot\mathcal{D}_{n}(\theta)\}\mathrm{d}\theta}.$
Remark 1.
Exact Bayesian inference is recovered from (1) by taking $\omega=1$ and setting $\mathcal{D}_{n}(\theta)=-\log p_{\theta}^{(n)}(y)$, where $p_{\theta}^{(n)}(y)$ is the assumed model density. Further, if the loss is additive, so that $\mathcal{D}_{n}(\theta)=\sum_{i=1}^{d}d(y_{i},\theta)$ for some known loss function $d:\mathcal{Y}\times\Theta\rightarrow\mathbb{R}_{+}$, we obtain the generalized posterior of Bissiri et al., (2016). The measure $\pi(\theta\mid\mathcal{D}_{n})$ is often referred to as a Gibbs posterior, but it is sometimes also called a generalized posterior.
Gibbs posteriors implicitly target the value of $\theta\in\Theta$ that minimizes the (limiting) expected loss, $\theta_{\star}:=\operatornamewithlimits{argmin\,}_{\theta\in\Theta}\lim_{n}\mathbb{E}n^{-1}\mathcal{D}_{n}(\theta)$, where $\mathbb{E}(\cdot)$ denotes expectation under $P^{(n)}_{0}$, and it is this quantity onto which they produce inferences. Unfortunately, the posterior $\pi(\theta\mid\mathcal{D}_{n})$, does not deliver inferences on $\theta_{\star}$ that are ‘calibrated’: a credible set for $\theta_{\star}$ based on $\pi(\theta\mid\mathcal{D}_{n})$ and having posterior probability $(1-\alpha)$ is calibrated if the credible set asymptotically contains $\theta_{\star}$ with $P^{(n)}_{0}$ probability $(1-\alpha)$. That is, a $(1-\alpha)$ credible set for $\pi(\theta\mid\mathcal{D}_{n})$ does not have coverage $(1-\alpha)$ under $P^{(n)}_{0}$.
The lack of calibration is due to the fact that, asymptotically, the Gibbs posterior quantifies uncertainty using an incorrect covariance matrix. In large samples, the posterior $\pi(\theta\mid\mathcal{D}_{n})$ behaves like the Gaussian density with mean $\theta_{\star}$, and variance $[\omega n\mathcal{H}(\theta_{\star})]$, i.e., like $N\{\theta;\mu=\theta_{\star},V=[\omega n\mathcal{H}(\theta_{\star})]\}$, where $\mathcal{H}_{n}(\theta):=n^{-1}\nabla_{\theta\theta}^{2}\mathcal{D}_{n}(\theta)$ denote the Hessian of $n^{-1}\mathcal{D}_{n}(\theta)$, and $\mathcal{H}(\theta):=\lim_{n}\mathbb{E}\{\mathcal{H}_{n}(\theta)\}$ its limiting value. Conversely, the posterior mean of $\pi(\theta\mid\mathcal{D}_{n})$ will converge towards $\theta_{\star}$ but have an asymptotic variance that is of the sandwich form: $\Sigma_{\star}:=\mathcal{H}(\theta_{\star})^{-1}\mathcal{I}(\theta_{\star})\mathcal{H}(\theta_{\star})^{-1}/n$. Thus, credible sets for $\theta_{\star}$ with posterior probability $(1-\alpha)$ do not in general contain $\theta_{\star}$ with $P^{(n)}_{0}$ probability $(1-\alpha)$.
Hence, posterior credible sets built from Gibbs posteriors are not calibrated in general, and Bayesian uncertainty quantification may not be reliable in many empirical applications. The lack of calibrated Bayesian inference in misspecified model has led researchers to consider approaches that ‘correct’ this issue. However, the suggested approaches of which we are aware are either based on computationally onerous bootstrapping procedures, which must bootstrap the entire posterior distribution multiple times and necessitates running many MCMC samplers - one for each bootstrap replication of the data - as well as ensuring that each MCMC sampler converges to the appropriate distribution, such as in Huggins and Miller, (2019), or Matsubara et al., (2022). Alternatively, one could apply ex-post corrections to $\pi(\theta\mid\mathcal{D}_{n})$ that require an explicit Gaussianity assumption on $\pi(\theta\mid\mathcal{D}_{n})$ to deliver calibrated inference; see Section 5.2 for further discussion on such approaches.
2.2 Reliable Bayesian Uncertainty Quantification
The lack of reliable Bayesian uncertainty quantification is a direct consequence of the mismatch between posterior variability, determined by $\mathcal{H}(\theta_{\star})^{-1}$, and the variability of the posterior mean, determined by $\Sigma^{-1}$. In principle, we would prefer to obtain a Gibbs posterior that does not meaningfully differ from the original posterior $\pi(\theta\mid\mathcal{D}_{n})$, and which also correctly quantifies uncertainty. It turns out that to produce such a posterior, we must only replace the original loss function $\mathcal{D}_{n}(\theta)$ by
$\mathcal{Q}_{n}(\theta):=\frac{1}{2}\log|W_{n}(\theta)|+n\cdot Q_{n}(\theta),\;\text{ where }Q_{n}(\theta):=\frac{1}{2}\cdot{\overline{m}_{n}(\theta)^{\top}}{}W_{n}(\theta)^{-1}{\overline{m}_{n}(\theta)}{}.$
$\overline{m}_{n}(\theta)=\nabla_{\theta}\mathcal{D}_{n}(\theta)/n$ and $W_{n}(\theta)$ is a $(d_{\theta}\times d_{\theta})$-dimensional covariance matrix. Replacing $\mathcal{D}_{n}(\theta)$ in equation (1) wtih $\mathcal{Q}_{n}(\theta)$ produces the Gibbs posterior
$\pi(\theta\mid\mathcal{Q}_{n}):=\frac{M_{n}(\theta)^{-1/2}\exp\{-n\cdot Q_{n}(\theta)\}\pi(\theta)}{\int_{\Theta}M_{n}(\theta)^{-1/2}\exp\{-n\cdot Q_{n}(\theta)\}\pi(\theta)\mathrm{d}\theta},$ (2)
where $M_{n}(\theta)=|W_{n}(\theta)|$, and where the notation $\pi(\theta\mid\mathcal{Q}_{n})$ encodes the posterior’s dependence on $Q_{n}(\theta)$. Unlike standard Gibbs posteriors, $\pi(\theta\mid\mathcal{Q}_{n})$ does not require a learning rate that must be tuned to ensure reasonable posterior coverage, and is equivalent to setting $\omega$ in (1) to unity.
The posterior $\pi(\theta\mid\mathcal{Q}_{n})$ is based on a weighted random quadratic-form in the “scores equations” $\overline{m}_{n}(\theta)$. Hence, throughout the remainder we refer to $\pi(\theta\mid\mathcal{Q}_{n})$ as the Q-posterior. We also clarify that while $Q_{n}(\theta)$ resembles a quadratic approximation, this does not imply that $\pi(\theta\mid\mathcal{Q}_{n})$ resembles a Gaussian distribution. More generally, $\pi(\theta\mid\mathcal{Q}_{n})$ remains meaningful when the parameters are defined over a restricted space, such as when the posterior places mass near the boundary of its support, or when the original posterior is multi-modal.
To intuitively explain why $\pi(\theta\mid\mathcal{Q}_{n})$ correctly quantifies uncertainty, consider a quadratic expansion of $Q_{n}(\theta)$ around $\theta_{n}=\arg_{\theta\in\Theta}\{0=\overline{m}_{n}(\theta)\}$, which, under regularity conditions, yields
$Q_{n}(\theta)-Q_{n}(\theta_{n})=-\frac{1}{2}(\theta-\theta_{n})^{\top}\Delta_{n}(\theta_{n})(\theta-\theta_{n})+R_{n}(\theta),\quad\Delta_{n}(\theta):=\mathcal{H}_{n}(\theta)^{\top}W_{n}(\theta)^{-1}\mathcal{H}_{n}(\theta)$
for some remainder term $R_{n}(\theta)$ whose order can be suitably controlled. Consequently, in large samples the Q-posterior behaves like
$\pi(\theta\mid\mathcal{Q}_{n})\propto\exp\left\{-\frac{n}{2}(\theta-\theta_{n})^{\top}\Delta_{n}(\theta_{n})(\theta-\theta_{n})+R_{n}(\theta)\right\}\pi(\theta).$
Since $\Delta_{n}(\theta_{n})$ is of the sandwich form and the above posterior kernel is sufficient to ensure correct uncertainty quantification (at least asymptotically).
Remark 2.
The above discussion suggests that $W_{n}(\theta)$ must be a consistent estimator for the variance of $\overline{m}_{n}(\theta_{n})$ in order for the Q-posterior to deliver calibrated inferences. Critically, however, specification of $W_{n}(\theta)$ does not require performing inference on any additional parameters. Further, $W_{n}(\theta)$ can often be taken as the sample variance of $\overline{m}_{n}(\theta_{\star})$ or some robust versions thereof if one suspects that the sample variance will not be consistent.
Remark 3.
Since the Q-posterior depends on $W_{n}(\theta)^{-1}$, calculation of this matrix may be costly when $d_{\theta}$ is large and may lead to singularity issues. In such situations, a recursive approximation to $W_{n}^{-1}(\theta)$ could instead be used. Focus on the case where $\overline{m}_{n}(\theta)=\sum_{i=1}^{n}m_{i}(\theta)/n$, $W_{n}(\theta)$ is the sample variance and define $\phi_{i}=\{m_{i}(\theta)-\overline{m}_{n}(\theta)\}/\sqrt{n}$. Define $\Psi_{0}=\varepsilon I_{d_{\theta}}$ for some known $\varepsilon>0$. By the Sherman Morrison formal, the inverse of the matrix
$\Psi_{n}(\theta)=\Psi_{0}+W_{n}(\theta)=\Psi_{0}+\sum_{i=1}^{n}\phi_{i}(\theta)\phi_{i}(\theta)^{\top}$
is given by
$\Psi_{n}^{-1}(\theta)=\Psi^{-1}_{n-1}{}(\theta)-\{1+\phi_{n}(\theta)^{\top}\Psi^{-1}_{n}(\theta){}\phi_{n}(\theta)\}^{-1}\Psi^{-1}_{n-1}(\theta){}\phi_{n}(\theta)\phi_{n}(\theta)^{\top}\Psi^{-1}_{n-1}(\theta){}.$
Following Cénac et al., (2025), we can take as an approximation to $\Psi_{n}(\theta)$ the matrix
$\Psi_{n}^{-1}(\theta)=\Psi^{-1}_{n-1}(\theta)-\{1+\phi_{n}(\theta)^{\top}\Psi_{n-1}^{-1}(\theta)\phi_{n}(\theta)\}^{-1}\Psi^{-1}_{n-1}(\theta)\phi_{n}(\theta)\phi_{n}(\theta)^{\top}\Psi^{-1}_{n-1}(\theta),$
which can be calculated recursively starting at $\Psi_{0}^{-1}=\varepsilon^{-1}I_{d_{\theta}}$ and does not require inverting any matrices. This approach has been used in the context of second-order optimization methods by Cénac et al., (2025) and in variational Bayes by Godichon-Baggioni et al., (2024), where the latter show that $\Psi_{n}$ converges to $\mathcal{I}(\theta)$ in large samples under various regularity conditions.
Remark 4.
The restriction of our analysis to fixed $d_{\theta}$ rules out the case of high-dimensional models, and limits the analysis conducted herein. However, due to the lack of a generative modeling structure, existing applications of Gibbs posteriors are often limited to cases where $d_{\theta}$ is not too large. Given this, we focus on the case of low-to-moderate $d_{\theta}$ and leave the extension of our analysis to cases where $d_{\theta}$ can increase with $n$ for future research. That being said, we remark that in the experiments that follow, we have applied the Q-posterior in problems where $d_{\theta}$ is up to twenty, using simple sample variance estimators and received reliable results with reasonable computing times.
Before formally demonstrating that $\pi(\theta\mid\mathcal{Q}_{n})$ correctly quantifies uncertainty, we present several examples which empirically illustrate that the Q-posterior delivers reliable uncertainty quantification across different choices of loss functions, and assumed models. We also show in Section A.5 of the supplementary material that the Q-posterior has a closed form when the assumed model is in the exponential family.
3 Examples
We now apply the Q-posterior approach to several examples that have been considered in the literature on standard and generalized Bayesian inference, including two examples of doubly intractable distributions.
3.1 Linear Regression
Consider the standard linear regression model
$y_{i}=x_{i}^{\top}\beta+\epsilon_{i},\;(i=1,\dots,n),$
where the error distribution is assumed to be $\epsilon_{i}\stackrel{{\scriptstyle iid}}{{\sim}}N(0,\sigma^{2})$, with $\sigma>0$ unknown, and $x_{i}$ and $\beta$ are $3\times 1$-dimensional vectors, where $x_{i,1}=1$ for all $i=1,\dots,n$. For simplicity, we consider flat priors on $\beta$ and $\sigma$. While the mean component of the regression is correctly specified, the true error term is given by $\epsilon_{i}\stackrel{{\scriptstyle iid}}{{\sim}}N(0,1/3+1/3\cdot|x_{2,i}|^{\gamma}+1/3\cdot|x_{3,i}|^{\gamma}),$ where $\gamma\geq 0$, and $x_{j,i}$ denotes the $j$th element of the vector $x_{i}$. When $\gamma=0$ the assumed model with homoskedastic errors is correctly specified, whereas if $\gamma\neq 0$, the assumed model is misspecified. Under correct specification, i.e. $\gamma=0$, the true value of the standard deviation of the noise under the assumed model is $\sigma=1$.
The goal of this example is to compare the accuracy of the standarad posterior and Q-posterior in correctly specified and misspecified regimes. As a competitor we also consider the infeasible method that correctly models the form of the heteroskedasticity parametrically. In particular, we consider posterior inference based on the assumed model
$y_{i}=x_{i}^{\top}\beta+\epsilon_{i},\quad\epsilon_{i}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}N\left(0,\xi_{1}+\xi_{2}|x_{2,i}|^{\gamma}+\xi_{3}|x_{i,3}|^{\gamma}\right),\quad(i=1,\dots,n),$ (3)
where $\xi_{1},\xi_{2},\xi_{3}\geq 0$ are unknown coefficients, each with flat priors, and $\gamma$ is known. Such an approach is infeasible since the form of heteroskedasticity is unknown in practice; however, since the form is known in this context, we consider the posterior for $\beta$ obtained from the model in (3) as an oracle benchmark. Under this heteroskedastic robust Bayes (HrB) approach we assume that $\gamma$ is set to its true value.
We generate $n=100$ observations from the model under $\gamma=0$, and $\gamma=2$; $x_{1,i}=1$ for all $i$, so that $\beta_{1}$ is the intercept, and $(x_{2,i},x_{3,i})^{\top}$ is generated as bivariate independent standard Gaussian (for each replication), and we set the true values as $\beta=(1,1,1)^{\top}$. When the model is misspecified in this manner, the pseudo-true remains equal to the true value of $\beta$, while the pseudo-true value of $\sigma$ is no longer unity, and so we focus in this example on inferences for $\beta$. We replicate this design 1000 times to create 1000 observed datasets of size $n=100$. For each dataset we sample the exact posterior, Q-posterior, and HrB-posterior using random walk Metropolis-Hastings (RWMH) with a Gaussian proposal kernel using the AdaptiveMCMC.jl library (Vihola, , 2014) in Julia. All posteriors are approximated using 20000 samples with an initial 1000 iterations for burn-in. For each replication, the matrix $W_{n}(\theta)$ in the Q-posterior is the sample variance of the score equations (based on the data in that replication). Across each dataset and method, we compare the posterior bias, variance and marginal coverage for the regression coefficients.
Table 1 summarizes the results, and shows that the Q-posterior produces results that are similarly located to exact Bayes, but has larger posterior variance under both regimes. In the homoskedastic regime ($\gamma=0$), the exact posterior and the Q-posterior both have coverage rates around 95%. In the heteroskedastic regime ($\gamma=2$), the coverage of exact Bayes is further away from the nominal level than the Q-posterior, with the lowest level of coverage around 87%.
The table also shows that the coverage of the HrB-posterior and the Q-posterior are similar, with the HrB-posterior having tighter credible intervals due to the fact that it correctly models the heteroskedasticity, whereas the Q-posterior does not directly model this quantity. Critically, however, the Q-posterior obtains similar results to the infeasible HrB-posterior without modelling the heteroskedastic variance. This feature is extremely useful in practice since the only way to reliably model heteroskedasticity is to use nonparametric Bayesian methods, which injects a substantial level of complexity into an otherwise simple inference problem.
Q-posterior Exact Bayes HrB
$\gamma=0$ Bias Var Cover Bias Var Cover Bias Var Cover
$\beta_{1}$0.0064 0.0109 0.947 -0.0067 0.0105 0.946 -0.0067 0.011 0.948
$\beta_{2}$0.001 0.0124 0.959 -0.0009 0.0107 0.948 -0.0008 0.0113 0.956
$\beta_{3}$0.0002 0.0125 0.963 -0.0004 0.0107 0.954 -0.0005 0.0112 0.958
$\gamma=2$ Bias Var Cover Bias Var Cover Bias Var Cover
$\beta_{1}$0.0068 0.011 0.956 -0.0074 0.0104 0.942 -0.0048 0.008 0.947
$\beta_{2}$0.0008 0.0201 0.963 -0.0004 0.0105 0.872 -0.0011 0.0139 0.952
$\beta_{3}$0.001 0.0204 0.965 -0.0015 0.0105 0.871 0.0002 0.0139 0.959
Table 1: Results in the normal linear regression model for exact Bayes, HrB-posterior and Q-posterior. Bias is the bias of the posterior mean across the replications. Var is the average posterior variance deviation across the replications. Cover is the actual posterior coverage, where the nominal level is set to 95% for the experiments.
We also explore how the Q-posterior scales to higher dimensional problems. We consider the same DGP but with $d=20$ regressors and $n=1000$. In the DGP $x_{4}$$x_{20}$ do not appear in the model of the variance for the DGP, and all other regression coefficients are set to zero. We consider the correctly specified ($\gamma=0$) and the misspecified ($\gamma=2$) cases. Here, the HrB model is defined as
$y_{i}=x_{i}^{\top}\beta+\epsilon_{i},\quad\epsilon_{i}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}N\left(0,\sum_{k=1}^{20}\xi_{k}|x_{k,i}|^{\gamma}\right),\quad(i=1,\dots,n).$
As the model is higher dimensional, we instead use 50000 MCMC iterations after 1000 burnin iterations, and we repeat the analysis for 100 independent datasets. The results are shown in Tables 2 for three specific covariates, and the results for the entire vector of covariates can be found in Appendix A.1.1 in the supplementary material. In the correctly specified case, the Q-posterior delivers inferences that are close to the exact. The HrB model produces average variances that are significantly larger, and in general there is some minor over-coverage. The inflated variances can be attributed to having to estimate twice the number of parameters where many of the true $\xi$ parameter values are 0. For the misspecified model, exact inference results in undercoverage for $\beta_{2}$ and $\beta_{3}$, whereas the Q-posterior delivers accurate coverage. The results for the Q-posterior are similar to the exact posterior for other parameters. Again the average posterior variances are larger for HrB, which exhibits some overcoverage.
Q-posterior Exact Bayes HrB
$\gamma=0$ Bias Var Cover Bias Var Cover Bias Var Cover
$\beta_{2}$0.0012 0.001 0.97 -0.0009 0.001 0.96 -0.0023 0.0017 0.98
$\beta_{3}$0.002 0.001 0.95 -0.0019 0.001 0.95 -0.0019 0.0017 0.96
$\beta_{20}$ 0.0035 0.001 0.95 0.0035 0.001 0.95 0.0037 0.0018 0.99
$\gamma=2$ Bias Var Cover Bias Var Cover Bias Var Cover
$\beta_{2}$0.0024 0.0017 0.95 -0.002 0.001 0.87 0.0002 0.002 0.96
$\beta_{3}$0.0029 0.0017 0.95 -0.0026 0.001 0.85 -0.0006 0.0019 0.92
$\beta_{20}$ 0.0044 0.001 0.98 0.0043 0.001 0.97 0.0039 0.0018 0.97
Table 2: Results in the normal linear regression model for exact Bayes, HrB-posterior and Q-posterior. Bias is the bias of the posterior mean across the replications. Var is the average posterior variance deviation across the replications. Cover is the actual posterior coverage, where the nominal level is set to 95% for the experiments.
3.2 Poisson Regression
For $i=1,\dots,n$ we observe count response data $y_{i}\in\mathbb{N}$, and covariates $x_{i}$, a $d_{\theta}\times 1$-dimensional vector with $x_{i,1}=1$, and our goal is inference on the unknown regression parameter $\theta$ in the generalized linear model (GLM):
$\mathbb{E}(y_{i}\mid x_{i})=\mu_{i}=g(x_{i}^{\top}\theta),\quad\text{var}(y_{i}\mid x_{i})=V(\mu_{i};\psi),$
where $g(\cdot)$ is a strictly monotone and differentiable link function, and $V(\cdot)$ is a positive and continuous variance function with dispersion parameter $\psi$. We have prior beliefs $\pi(\theta)\propto 1$, while we treat $\psi$ as a hyper-parameter. For observed counts a common choice is to model the distribution as Poisson with link function $g(\cdot)=\exp(\cdot)$.
A useful alternative to conducting standard Bayesian inference in this setting, is to instead consider generalized Bayesian using the quasi-likelihood of Wedderburn, (1974) for the Poisson model with variance function $V(\mu_{i};\psi)=\psi\mu_{i}$. Such an approach is equivalent to implementing (1) under the loss function
$\mathcal{D}_{n}(\theta)=-\psi^{-1}\sum_{i=1}^{n}\left\{y_{i}\log(\mu_{i})-\mu_{i}\right\},$
with $\psi>0$ attempting to account for over-dispersion. Applying this choice within $\mathcal{D}_{n}(\theta)$, and taking $\omega=1$, produces the Gibbs posterior
$\pi(\theta\mid\mathcal{D}_{n},\psi)\propto\pi(\theta)\exp\left[\psi^{-1}\sum_{i=1}^{n}\left\{y_{i}x_{i}^{\top}\theta-\exp(x_{i}^{\top}\theta)\right\}\right].$
This approach has been suggested in several studies, such as Ventura and Racugno, (2016), and has been shown by Agnoletto et al., (2023) to produce asymptotically correct levels of calibration when the true variance function takes the form $\text{var}(y_{i}\mid x_{i})=V(\mu_{i};\psi)$. Following Agnoletto et al., (2023), estimation of $\psi$ can be obtained by first fitting an over-dispersed Poisson GLM to obtain the point estimator $\hat{\psi}$, with MCMC then used to sample $\pi(\theta\mid\mathcal{D}_{n},\hat{\psi})$.
While over-dispersion is a common argument for using classes of count distributions other than the Poisson, specification of the dispersion functions is not particularly easy, and necessitates joint inference on $\theta$ and $\psi$, even though inference on $\theta$ is the goal. In this section, we compare standard Bayesian inference using the Poisson model, the approach of Agnoletto et al., (2023) based on $\pi(\theta\mid\mathcal{D}_{n},\hat{\psi})$, and the Q-posterior based on the assumed Poisson approximating model, which is equivalent to fixing $V(\mu_{i};\psi)=\mu_{i}$. [^2] Similar to the linear regression model, we take as our matrix $W_{n}(\theta)$ in the Q-posterior the sample covariance of the score equations. We follow the simulation design of Agnoletto et al., (2023) and conduct this comparison in the case of over-dispersed counts. Our results show that the Q-posterior delivers results that are just as reliable as the approach of Agnoletto et al., (2023) but without needing to model the dispersion function $V(\mu_{i};\psi)$ or conduct inference on the hyper-parameter $\psi$.
Following the simulation design of Agnoletto et al., (2023) the data is generated as follows: for $\psi_{0}$ some true dispersion parameter (fixed at $\psi_{0}=1.5$ in our experiments), first generate $\tilde{y}_{i}\stackrel{{\scriptstyle iid}}{{\sim}}\text{Ga}\{\mu_{i}/\psi_{0},1/\psi_{0}\}$, and define the observed $y_{i}$ as the closest integer value of $\tilde{y}_{i}$. This process yields counts that are over-dispersed, so that the Poisson model is overly-precise. We again take $x_{i}$ and $\theta$ as a $d_{\theta}\times 1$-dimensional vector, where $x_{i,1}=1$ for all $i=1,\dots,n$, and $(x_{2,i},\ldots,x_{d_{\theta},i})^{\top}$ is generated as multivariate independent Gaussian. We consider two simulation designs corresponding to $d_{\theta}=10$ and $d_{\theta}=20$: when $d_{\theta}=10$ we set $\theta=(3.5,0.5,-0.5,0.5,0,\dots,0)^{\top}$, and for $d_{\theta}=20$ the parameter $\theta$ has the same structure but the remaining ten entries in $\theta$ are zero.
For $n=1000$ we generate 100 replicated datasets from this DGP, and present the bias of the posterior mean, the posterior variance and the marginal coverage of each method across the two regimens for $\gamma$ [^3] Posterior sampling for the generalized and Q-posteriors is carried out using STAN along with the default choices for the NUTS algorithm. We use two chains, a warmup period of 1000 iterations and 3000 iterations in total. Table 3 displays the results for the first four covariates in both cases and demonstrates that the Q-posterior and generalized Bayes posterior behave very similarly and have coverages that are close to the nominal level. In contrast, standard Bayesian inference based on the Poisson model is overly precise, and has much poorer coverage. Critically, unlike the approach of Agnoletto et al., (2023), the Q-posterior obtains reliable coverage without having to model the dispersion or estimate the over-dispersion parameter $\psi$.
Q-posterior Exact Bayes GenBayes
$d_{\theta}=10$ Bias Var Cover Bias Var Cover Bias Var Cover
$\theta_{1}$1.44 0.055 0.92 -0.915 0.037 0.88 -0.945 0.055 0.93
$\theta_{2}$ 1.36 0.032 0.96 1.319 0.022 0.94 1.317 0.032 0.97
$\theta_{3}$0.45 0.032 0.95 -0.407 0.021 0.88 -0.398 0.032 0.95
$\theta_{4}$0.16 0.032 0.95 -0.219 0.022 0.88 -0.222 0.032 0.98
$d_{\theta}=20$ Bias Var Cover Bias Var Cover Bias Var Cover
$\theta_{1}$1.44 0.055 0.92 -0.915 0.037 0.88 -0.945 0.055 0.93
$\theta_{2}$ 1.36 0.032 0.96 1.319 0.022 0.94 1.317 0.032 0.97
$\theta_{3}$0.45 0.032 0.95 -0.407 0.021 0.88 -0.398 0.032 0.95
$\theta_{4}$0.16 0.032 0.95 -0.219 0.022 0.88 -0.222 0.032 0.98
Table 3: Results in the Poisson regression model for exact Bayes, Q-posterior, and generalized Bayes (GenBayes) based on the quasi-likelihood. Bias is the bias of the posterior mean across the replications and has been multiplied by 100 for readability. Var is the average posterior variance deviation across the replications and has been multiplied by 100 for readability. Cover is the actual posterior coverage, where the nominal level is set to 95% for the experiments.
3.3 Doubly Intractable Models
In many interesting settings the likelihood for the model $P_{\theta}^{(n)}$ is of the form $p_{\theta}^{(n)}(y)=Z_{\theta}^{-1}\tilde{p}(y\mid\theta)$, where $\tilde{p}(y\mid\theta)$ is an analytically tractable density kernel, and $Z_{\theta}$ is an intractable normalizing constant. Bayesian inference in such settings, often called doubly intractable models since both $Z_{\theta}$ and the marginal likelihood are intractable, is challenging and often requires resorting to an approximation of $p_{\theta}^{(n)}(y)$. Examples include spatial models, exponential random graphs, and certain discrete data settings (Matsubara et al., , 2022).
Matsubara et al., (2022, 2023) demonstrate how generalized Bayesian methods based on certain classes of discrepancies can be used to deliver posterior inferences in such settings. In both cases, the key insight is that by replacing the likelihood with a well-chosen discrepancy function, computation of the intractable normalizing constant can be circumvented. For continuous variables, this choice leads Matsubara et al., (2022) to produce Bayesian inference using the kernel Stein discrepancy (KSD-Bayes); for discrete data, Matsubara et al., (2023) use the discrete Fisher divergence (DFD) to produce generalized Bayes posteriors (DFD-Bayes).
Regardless of which divergence is used to produce the generalized Bayesian posterior for $\theta$, the resulting posterior is not calibrated in general. To circumvent this issue, Matsubara et al., (2023) propose a computationally onerous bootstrapping procedure to deliver calibrated inferences, while Matsubara et al., (2022) propose an approximate calibration procedure based on a particular choice for the posterior learning rate.
Theorem 1 demonstrates that so long as $p_{\theta}^{(n)}$ is continuously differentiable, the Q-posterior based on the KSD (or DFD) delivers calibrated inferences without bootstrapping, or the need to choose the learning rate.
3.3.1 DFD-Bayes: Conway-Maxwell-Poisson Model
This section performs approximate Bayesian inference for the Conway-Maxwell-Poisson model (Conway and Maxwell, , 1962), which is a flexible model for discrete data $x$, with $x\geq 0$, that can capture both under- and over-dispersion. The probability mass function for a single observation $x$ conditional on parameter $\theta=(\theta_{1},\theta_{2})^{\top}$ is
$p(x\mid\theta)$ $=\frac{\tilde{p}(x|\theta)}{Z_{\theta}},$
where $\tilde{p}(x|\theta)=(\theta_{1})^{x}(x!)^{-\theta_{2}}$, with $\theta_{1}>0$ and $\theta_{2}\in[0,1]$. The normalising constant $Z_{\theta}=\sum_{y=0}^{\infty}\tilde{p}(y|\theta)$ does not have a closed form expression, except for special cases such as when $\theta_{2}=1$, which recovers the Poisson distribution with mean $\theta_{1}$. However, the normalising constant can be computed to high-accuracy with little computational effort, which facilitates comparison with a highly precise approximation of the true posterior.
To avoid the intractable normalising constant associated with discrete distributions, Matsubara et al., (2023) conduct generalized Bayesian inference on $\theta$ using the DFD as the loss function. In the one-dimensional data setting considered here the DFD between the statistical model conditioned on $\theta$, $p(\cdot|\theta)$ and the empirical distribution of the data, $p_{n}$, is defined as
$\mbox{DFD}\{p(\cdot|\theta)||p_{n}\}$ $=\frac{1}{n}\sum_{i=1}^{n}\left(\frac{p(x_{i}^{-}|\theta)}{p(x_{i}|\theta)}\right)^{2}-2\left(\frac{p(x_{i}|\theta)}{p(x_{i}^{+}|\theta)}\right),$
and it is evident that the intractable normalising constants cancel in the ratio. For this example, $x^{+}=x+1$ and $x^{-}=x-1$, unless $x=0$ in which case we set $x^{-}=\mathrm{max}\{y_{i}\}_{i=1}^{n}$, i.e. the maximum value of the dataset.
Matsubara et al., (2023) embed the DFD within a generalised Bayes framework to conduct approximate Bayesian inferences without invoking the $\theta$-dependent normalisation constant. To calibrate the scaling parameter $\omega$ in the generalized posterior, Matsubara et al., (2023) propose the following steps: first, $B$ bootstrap replications of the observed data are created, and for each of the bootstrapped datasets we obtain an estimator of $\theta$, say $\{\theta_{n}^{(b)}:b=1,\dots,B\}$, by minimizing the DFD between the assumed model, and the $b$-th bootstrapped dataset; the value of $\omega$ is then chosen to minimize the Fisher divergence between the generalised posterior and the empirical bootstrap sample $\{\theta_{n}^{(b)}:b=1,\dots,B\}$. Once this value of $\omega$ is obtained, it can then be used within an MCMC algorithm to generate samples from the posterior.
In contrast, the Q-posterior approach avoids a calibration process, and should deliver calibrated Bayesian inferences. All that is required is to compute the score of each component of the DFD, which in this case can be done analytically. The implementation can be accelerated by computing the score for each unique value of the dataset, and weighting by the number of replicates of each unique value. Again, the AdaptiveMCMC.jl library in Julia is used to sample the Q-posterior.
Based on 100 independent datasets of size $n=2000$ using the true parameter value $\theta=(4,0.75)^{\top}$. Table 4 shows the results for the Q-posterior, which is compared with an accurate approximation of the true posterior and the GBI bootstrap approach of Matsubara et al., (2023). Although not formally proved in Matsubara et al., (2023), we conjecture that, provided the model is correctly specified, the pseudo-true parameter value, i.e. the one that coincides with the minimiser of the DFD as $n\rightarrow\infty$, is equal to the true parameter value. Indeed, we consider a dataset of size 10 million and find that the minimiser of the DFD coincides exactly with the true parameter when rounded to two decimal place accuracy. The Q-posterior approach produces posteriors with a larger standard deviation than the true posterior, but still achieves reasonable coverage rates. The Q-posterior results are slightly less accurate than the GBI approach of Matsubara et al., (2023), however, we remark that the approach of Matsubara et al., (2023) is specifically focused on intractable discrete models whereas the Q-posterior is generally applicable. Further, the Q-posterior avoids the need to bootstrap the distribution of the point estimator that minimises the loss.
Figure 1 displays posterior approximations for a single dataset. As is evident, the Q-posterior approach produces an approximation that is inflated relative to the true posterior, and is similar to the GBI results of Matsubara et al., (2023). We find that for some of the datasets, the Q-posterior has a substantially heavier tail (see Figure 2) than the GBI approach, which for the parameter $\theta_{1}$ seems to lead to some inflation of the MSE and bias shown in Table 4.
Method MSE Bias St. Dev. Coverage (90%)
$\theta_{1}$
True 0.11 0.076 0.24 94
Q-posterior 0.60 0.27 0.51 85
Matsubara et al., (2023) 0.28 0.14 0.40 94
$\theta_{2}$
True 0.0039 0.012 0.045 93
Q-posterior 0.012 0.040 0.077 87
Matsubara et al., (2023) 0.0074 0.019 0.061 93
Table 4: Results for the Conway-Maxwell-Poisson example using 100 independent datasets simulated with true parameter value $\theta=(\theta_{1},\theta_{2})^{\top}=(4,0.75)^{\top}$. For a detailed description on the column entries, see the caption in Table 11.
Figure 1: Univariate densities estimates of approximations to the posterior distribution for a single dataset generated from a CMP model with true parameter value $\theta=(\theta_{1},\theta_{2})^{\top}=(4,0.75)^{\top}$. Shown are posterior approximations based on the Q-posterior (green), the GBI approach of Matsubara et al., (2022) (red) and an accurate approximation of the exact likelihood (blue).
Figure 2: Univariate posterior standard deviations of $\theta_{1}$ across 100 datasets generated from a CMP model with true parameter value $\theta=(\theta_{1},\theta_{2})^{\top}=(4,0.75)^{\top}$. Shown are results based on the Q-posterior (green), the GBI approach of Matsubara et al., (2022) (red) and an accurate approximation of the exact likelihood (blue).
3.3.2 KSD-Bayes: Contaminated Normal
While discrete Fisher divergence-based Bayes (DFD-Bayes) can be used to conduct posterior inference in models where the normalizing constant for the mass function is intractable, Matsubara et al., (2022) suggest using the kernel Stein discrepancy (KSD) within a generalized Bayes framework when the variables under observation are continuous. In particular, to avoid the intractable normalizing constant $Z_{\theta}$ in $p^{(n)}_{\theta}(y)=p(y\mid\theta)/Z_{\theta}$, Matsubara et al., (2022) conduct generalized Bayesian inference using the KDS, which is defined as follows: for some positive-definite kernel function $K:\mathcal{Y}\times\mathcal{Y}\rightarrow\mathbb{R}^{d\times d}$, $\mathbb{P},\mathbb{Q}\in\mathcal{P}(\mathcal{Y})$, where $\mathcal{P}(\mathcal{Y})$ denotes the space of distributions over $\mathcal{Y}$, and $\mathcal{S}_{\mathbb{Q}}$ the Stein operator (see Matsubara et al., , 2022 for a formal definition), the kernel Stein discrepancy takes the form $\operatorname{KSD}^{2}(\mathbb{Q}\|\mathbb{P}):=\mathbb{E}_{Y,Y^{\prime}\sim\mathbb{P}}\left[\mathcal{S}_{\mathbb{Q}}\mathcal{S}_{\mathbb{Q}}K\left(Y,Y^{\prime}\right)\right],$ which can be explicitly estimated using
$\mathrm{KSD}^{2}({P}^{(n)}_{\theta}\|\mathbb{P}_{n})=\frac{1}{n^{2}}\sum_{i=1}^{n}\sum_{j=1}^{n}\mathcal{S}_{\mathbb{P}_{\theta}}\mathcal{S}_{\mathbb{P}_{\theta}}K\left(y_{i},y_{j}\right),$
where $\mathbb{P}_{n}$ is the empirical measure of the sample $y_{1},\dots,y_{n}$. Matsubara et al., (2022) show that $\mathrm{KSD}^{2}\left(\mathbb{P}_{\theta}\|\mathbb{P}_{n}\right)$ can be evaluated in closed form when $\mathcal{Y}=\mathbb{R}^{d}$ and when $\mathcal{S}_{\mathcal{Q}}$ is the Langevin-Stein operator; see Matsubara et al., , 2022 for complete details, and, in particular, the discussion surrounding their equation (5).
Matsubara et al., (2022) embed the KSD within a generalized Bayesian framework, and demonstrate that in exponential family models with conjugate priors, the KSD-Bayes posterior has a closed form. The main motivation behind the use of KSD-Bayes comes from its robustness properties, which allows the authors to show that the KSD-Bayes posterior displays a global bias-robustness property under data contamination.
While KSD-Bayes displays meaningful robustness, the resulting posterior is not calibrated. The authors suggest a two-stage approach for setting the learning rate to overcome this issue, however, the resulting KSD-Bayes posterior will not be calibrated unless $\theta$ is a scalar. In contrast, the Q-posterior based on $\mathrm{KSD}^{2}({P}^{(n)}_{\theta}\|\mathbb{P}_{n})$ inherits the bias robustness properties of the KSD-Bayes posterior, and also delivers calibrated inference.
To demonstrate this, we compare the behavior of the Q-posterior based on KSD loss and the KSD-Bayes in the contaminated normal location model analyzed in Section 6.1 of Matsubara et al., (2022). For both the Q-posterior and KSD-Bayes, we set the kernel function $K(\cdot,\cdot)$ as the inverse multi-quadratic kernel suggested as the default choice by Matsubara et al., (2022), and for KSD-Bayes we use their optimal choice of learning rate in an attempt to correct the coverage of the KSD-Bayes posterior; see Section 5 of Matsubara et al., (2022) for details on these choices.
Using the contaminated normal model design described in Matsubara et al., (2022), the assumed model is $P_{\theta}=\mathcal{N}(\theta,1)$, with prior belief $\theta\sim\mathcal{N}(0,1)$, but the actual generating process is an $\epsilon$-contamination of $P_{\theta}$ with contaminating distribution $\mathcal{N}(5,3)$. We take the true value of $\theta=0$, so that the true DGP is $(1-\epsilon)\mathcal{N}(0,1)+\epsilon\mathcal{N}(5,3)$. We generate 100 replicated datasets from the above DGP under $\epsilon=0$ (correct specification), and $\epsilon=0.10$ (model misspecification). For each dataset, we apply the Q-posterior and KSD-Bayes, and across the replications we calculate the MSE, bias and average posterior standard deviation, as well as the Monte Carlo coverage. Table 5 demonstrates that the Q-posterior has almost the exact same location as the KSD-Bayes posterior, but its credible sets are wider, and much closer to the nominal level than the KSD-Bayes posterior, even though the KSD-Bayes posterior utilizes the two-stage learning rate described in Matsubara et al., (2022). Similar findings hold under both designs.
Method MSE Bias St. Dev. Coverage (95%)
$\theta$, $\epsilon=0$
KSD 0.0163 -0.0129 0.1031 88
Q-posterior 0.0168 -0.0132 0.1211 93
$\theta$, $\epsilon=0.10$
KSD 0.0109 0.0684 0.1044 89
Q-posterior 0.0147 0.0687 0.1212 95
Table 5: Results for the contaminated normal location model with 100 independent datasets simulated with true parameter value $\theta=0$. The case of $\epsilon=0$ corresponds to correct specification, while $\epsilon=0.10$ corresponds to 10% data contamination. For a detailed description on the column entries, see the caption in Table 11.
4 Theoretical Results
4.1 Assumptions
Using standard regularity conditions regarding the behavior of $\overline{m}_{n}(\theta)$ and $W_{n}(\theta)$ that are commonly encountered in frequentist inference, we formally demonstrate that $\pi(\theta\mid\mathcal{Q}_{n})$ has calibrated credible sets. We first define several notations that are used to simplify the statement of our results. For a positive sequence $a_{n}\rightarrow+\infty$ as $n\rightarrow+\infty$, we say that $X_{n}=o_{p}(a_{n}^{-1})$ if the sequence $a_{n}X_{n}$ converges to zero in probability, while the notation $X_{n}=O_{p}(a_{n}^{-1})$ denotes that $a_{n}X_{n}$ is bounded in probability. For a set $A\subseteq\mathbb{R}^{d}$, $\mathrm{Int}(A)$ denotes the interior of $A$. The notation $\Rightarrow$ denotes weak convergence under $P^{(n)}_{0}$.
Assumption 1.
There exists a function $m:\Theta\rightarrow\mathbb{R}^{d_{\theta}}$ satisfying the following conditions.
  1. (i)
    $\sup_{\theta\in\Theta}\|\overline{m}_{n}(\theta)-m(\theta)\|=o_{p}(1)$.
  2. (ii)
    There exists $\theta_{\star}\in\mathrm{Int}(\Theta)$, such that $m(\theta_{\star})=0$.
  3. (iii)
    For some $\delta>0$, $m(\theta)$ is continuously differentiable over $\|\theta-\theta_{\star}\|\leq\delta$, and $\mathcal{H}(\theta_{\star}):=-\nabla_{\theta}m(\theta_{\star})$ is invertible.
  4. (iv)
    There is a positive-definite matrix $\mathcal{I}(\theta_{\star})$, such that $\sqrt{n}\overline{m}_{n}(\theta_{\star})\Rightarrow N\{0,\mathcal{I}(\theta_{\star})\}$.
  5. (v)
    For any $\epsilon>0$, there is a $\delta>0$ such that
    $\limsup_{n\rightarrow\infty}P^{(n)}_{0}\left(\sup_{\left\|\theta-\theta_{0}\right\|<\delta}\frac{\sqrt{n}\left\|\{\overline{m}_{n}(\theta)-\overline{m}_{n}(\theta_{\star})\}-\{m(\theta)-m(\theta_{\star})\}\right\|}{1+\sqrt{n}\|\theta-\theta_{\star}\|}>\epsilon\right)<\epsilon.$
Assumption 2.
The following conditions are satisfied for some $\delta>0$: (i) for $n$ large enough, the matrix ${W}_{n}(\theta)$ is positive semi-definite and symmetric for all $\theta\in\Theta$, and positive-definite for all $\|\theta-\theta_{\star}\|\leq\delta$; (ii) there exists a matrix $W(\theta)$, positive semi-definite, and symmetric, for all $\theta\in\Theta$, and such that $\sup_{\theta\in\Theta}\|{W}_{n}(\theta)-W(\theta)\|=o_{p}(1)$, and, for all $\|\theta-\theta_{\star}\|\leq\delta$, $W(\theta)$ is continuous and positive-definite; (iii) for any $\epsilon>0$, $\sup_{\|\theta-\theta_{\star}\|\geq\epsilon}m(\theta)^{\top}W(\theta)^{-1}m(\theta)>0$.
Remark 5.
Assumptions 1 and 2 together enforce smoothness and identification conditions on $Q_{n}(\theta)$; Assumption 1 is also similar to conditions maintained when studying the theoretical behavior of generalized posteriors (see, in particular, Lemmas 1 and 2 in Chernozhukov and Hong, , 2003). These conditions permit existence of a quadratic expansion of $Q_{n}(\theta)-Q_{n}(\theta_{\star})$ that is smooth in $\theta$ near $\theta_{\star}$, and with a remainder term that can be suitably controlled. Assumption 1.(i) is a uniform law of large numbers and is satisfied for a large class of functions under many different data generating processes, and can often be verified using method for empirical processes (see, e.g., Vaart and Wellner, , 2023, Chapter 3). Assumption 1.(v) is a smoothness condition on the score that does not require that $\overline{m}_{n}(\theta)$ is differentiable, but only that $\overline{m}_{n}(\theta)$ is asymptotically equicontinuous; Assumption 1.(iii) requires that the limit counterpart $m(\theta)$ is differentiable near $\theta_{\star}$. The weaker smoothness condition in Assumption 1.(v) is important as many loss functions are only weakly differentiable, e.g., loss functions based on absolute value functions, which satisfy Assumption 1.(v) but do not satisfy a standard differentiability condition. Assumption 1.(v) is often used in frequentist point estimation theory. We also note that Assumption 1 does not require that $\theta_{\star}$ is unique. This is important as we will later treat the case of multi-modal posteriors caused by score equations that admit multiple roots.
Remark 6.
Since the behavior of the Q-posterior depends on the matrix $W_{n}(\theta)$, we require certain regularity conditions on $W_{n}(\theta)$. Assumption 2(i) requires that, for $n$ large enough, the matrix $W_{n}(\theta)$ is positive-definite for any $\theta$ sufficiently close to $\theta_{\star}$, while Assumption 2(ii) requires that $W_{n}(\theta)$ converges uniformly to $W(\theta)$, which is continuous and positive-definite for all $\theta$ sufficiently close to $\theta_{0}$. Note that Assumption 2(ii) does not require $W(\theta)$ to be invertible uniformly over $\Theta$, but only sufficiently close to $\theta_{\star}$. This implies that the limiting quadratic form $Q(\theta)=2^{-1}m(\theta)^{\top}W(\theta)^{-1}m(\theta)$ need not be continuous over $\Theta$. Therefore, it is necessary to maintain Assumption 2(iii) to ensure identification. However, if $W(\theta)^{-1}$ is positive-definite over $\Theta$ this identification assumption is automatically satisfied.
The next assumption requires the existence of certain prior moments.
Assumption 3.
For $\theta_{\star}$ as defined in Assumption 1, $\pi(\theta_{\star})>0$ and $\pi(\theta)$ is continuous on $\Theta$. For some $p\geq 1$, $\int_{\Theta}\|\theta\|^{p}\pi(\theta)\mathrm{d}\theta<+\infty$, and for all $n$ large enough $\int_{\Theta}|W_{n}(\theta)|^{-1/2}\|\theta-\theta_{\star}\|^{p}\pi\left(\theta\right)\mathrm{d}\theta<+\infty$.
Remark 7.
The condition $\int_{\Theta}|W_{n}(\theta)|^{-1/2}\|\theta-\theta_{\star}\|^{p}\pi\left(\theta\right)\mathrm{d}\theta<+\infty$ is not commonly encountered and is required since the matrix $W_{n}(\theta)$ may be singular far away from $\theta_{\star}$. This condition requires that these singularities are controllable under the prior, and ensures that the posterior will be well-behaved even though $W_{n}(\theta)$ can have points of singularity. Similar to Assumptions 2 (iii), if $W_{n}(\theta)$ is invertible over $\Theta$, this condition is automatically satisfied.
Remark 8.
Since the vast majority of situations where Gibbs posteriors are applied are of low to moderate dimension, we have chosen to focus on the case where $d_{\theta}$ is fixed and of low-to-moderate dimension. In cases where $d_{\theta}$ is not fixed, or where the dimension is large relative to $n$, the above assumptions will not be valid and alternative arguments would be necessary to understand the behavior of the Q-posterior. Further, in such cases, one may have to apply alternative approximation techniques to produce posterior samples, as well as the matrix inverse approximation suggested in Remark 3 While interesting, we leave this extensions for future study.
4.2 Results
4.2.1 Unique identification
If the value $\theta_{\star}$ in Assumption 1 is unique, then the Q-posterior correctly quantifies uncertainty. To state this result simply, define $\vartheta:=\sqrt{n}(\theta-\theta_{n})$, $\Delta(\theta):=\mathcal{H}(\theta)W(\theta)^{-1}\mathcal{H}(\theta)$ and $\mathcal{T}_{n}:=\{\vartheta=\sqrt{n}(\theta-\theta_{n}):\theta\in\Theta\}$.
Theorem 1.
If Assumptions 1-3 hold, and $\theta_{\star}$ unique, then for $p=0,1$ $n\rightarrow\infty$,
$\int_{\mathcal{T}_{n}}\|\vartheta\|^{p}|\pi(\vartheta\mid\mathcal{Q}_{n})-N\{\vartheta;0,\Delta(\theta_{\star})^{-1}\}|\mathrm{d}\vartheta=o_{p}(1).$
Corollary 1.
Define $\overline{\theta}=\int_{\Theta}\theta\pi(\theta\mid\mathcal{Q}_{n})\mathrm{d}\theta$. If the assumptions of Theorem 1 hold, then
$\sqrt{n}(\overline{\theta}-\theta_{\star})\Rightarrow N\{0,\Delta(\theta_{\star})^{-1}\mathcal{H}_{\star}W(\theta_{\star})^{-1}\mathcal{I}_{\star}W(\theta_{\star})^{-1}\mathcal{H}_{\star}\Delta(\theta_{\star})^{-1}\}.$
Remark 9.
Theorem 1 shows that when $\theta_{\star}$ is unique the Q-posterior converges to a normal density with mean zero and variance $\Delta(\theta^{\star})^{-1}$ ($p=0$ in Theorem 1). The uniqueness of $\theta_{\star}$ the zero of the limit score equations $0=m(\theta)$ is necessary to ensure that the posterior shape is asymptotically Gaussian. Section 4.2.2 generalizes this to the case of multiple solutions. Uniqueness of $\theta_{\star}$ is a standard condition in the theoretical analysis of Bayesian methods, and is satisfied in many, but not all, contexts.
Remark 10.
Corollary 1 shows that the posterior mean of the Q-posterior converges to an asymptotically mean-zero normal random variable. However, Theorem 1 implies that the widths of credible sets obtained from $\pi(\theta\mid\mathcal{Q}_{n})$ are governed by $\Delta(\theta_{\star})^{-1}$. Hence, Theorem 1 and Corollary 1 together imply that for $\pi(\theta\mid\mathcal{Q}_{n})$ to be calibrated, we require that
$W(\theta_{\star})=\lim_{n\rightarrow+\infty}\text{Cov}\left\{\sqrt{n}\overline{m}_{n}(\theta_{\star})\right\}=\mathcal{I}_{\star}.$
When $\overline{m}_{n}(\theta)$ takes the form $\overline{m}_{n}(\theta)=n^{-1}\sum_{i=1}^{n}m_{i}(\theta)$, for some known functions $m_{i}(\theta)$ depending only on the $i$-th sample unit, the matrix
${W}_{n}(\theta)=n^{-1}\sum_{i=1}^{n}\left\{m_{i}(\theta)-\overline{m}_{n}(\theta)\right\}\left\{m_{i}(\theta)-\overline{m}_{n}(\theta)\right\}^{\top}$ (4)
often provides a consistent estimator for $\mathcal{I}_{\star}$. However, if it is unknown whether $W_{n}(\theta)$ in (4) is consistent for $\mathcal{I}_{\star}$, we suggest using robust covariance estimators, which can deliver consistent estimators of $\mathcal{I}_{\star}$ under very general conditions.
4.2.2 Non-unique Identification
Since the model is possibly misspecified, the limit score equation $0=m(\theta)$ may not admit a unique solution. If the score equation $0=m(\theta)$ is satisfied at multiple points, but the set of solutions $\Theta_{\star}:=\{\theta\in\mathrm{Int}(\Theta):0=m(\theta)\}$ only contains a finite number of well-separated points, then a result similar to Theorem 1 will be satisfied at each element in $\Theta_{\star}$. To derive such a result, recall $\Delta(\theta):=\mathcal{H}(\theta)W(\theta)^{-1}\mathcal{H}(\theta)$, let $\lambda_{\text{min}}\{\Delta(\theta)\}$ be the smallest eigenvalue of $\Delta(\theta)$, and impose the following assumption on $\Theta_{\star}$.
The following result demonstrates that, unlike Gaussian posterior corrections/approximations, the Q-posterior remains a meaningful object even if the posterior is multi-modal.
Assumption 4.
The set $\Theta_{\star}$ is finite and for some $\delta>0$, at least one $\theta\in\Theta_{\star}$, $0<\delta<\lambda_{\text{min}}\{\Delta(\theta)\}\leq 1/\delta$.
Similarly to the case where $\theta_{\star}$ is unique, we must choose a centering sequence to obtain the posterior shape. In general, a useful choice for this centering sequence would be $\theta_{n}:=\arg_{\theta\in\Theta}\{0=\overline{m}_{n}(\theta)\}$. However, Assumption 1 by itself does not imply that $\theta_{n}$ is consistent for any $\theta_{\star}\in\Theta_{\star}$, which means that without further assumptions $\theta_{n}$ cannot be used as a centering sequence to determine the posterior shape. [^4] In general, a sufficient condition for $\theta_{n}$ to be consistent for $\theta_{\star}$ is to require continuous differentiability of $\overline{m}_{n}(\theta)$, which is not maintained in Assumption 1. While it is possible to show the existence and consistency of $\theta_{n}=\arg_{\theta\in\Theta}\{0=\overline{m}_{n}(\theta)\}$ under additional smoothness conditions on $\overline{m}_{n}(\theta)$, to maintain generality, we impose the following high-level condition.
Assumption 5.
For some $\theta_{\star}\in\Theta$, $\theta_{n}=\arg_{\theta\in\Theta}\{0=\overline{m}_{n}(\theta)\}$ satisfies $\overline{m}_{n}(\theta_{n})=0$, and $\|\theta_{n}-\theta_{\star}\|=o_{p}(1)$.
To present the behavior of $\pi(\theta\mid\mathcal{Q}_{n})$ over $\Theta_{\star}$, let $\theta_{n}=\{0=\arg_{\theta\in\Theta}\overline{m}_{n}(\theta)\}$, define $t:=\sqrt{n}(\theta-\theta_{n})$ as the local parameter, $\mathcal{T}_{n}:=\{t=\sqrt{n}(\theta-\theta_{n}):\theta\in\Theta\}$ the support of $t$, and let $\pi(t\mid\mathcal{Q}_{n})=\pi(\theta_{n}+t/\sqrt{n}\mid\mathcal{Q}_{n})/{n}^{d_{\theta}/2}$ be the posterior for $t$.
Theorem 2.
If Assumptions 1-5 are satisfied, then for any finite $\gamma>0$, and some density function $\pi_{\star}(t)\propto N\{t;0,\Delta(\theta_{\star})^{-1}\}$, as $n\rightarrow+\infty$,
$\left|\int_{\|t\|\leq\gamma}{\pi}(t\mid\mathcal{Q}_{n})\mathrm{d}t-\int_{\|t\|\leq\gamma}\pi_{\star}(t)\mathrm{d}t\right|=o_{p}(1).$
Theorem 2 demonstrates that if the score equations $0=m(\theta)$ admit multiple roots, say $\theta_{1,\star}$ and $\theta_{2,\star}$, then the Q-posterior resembles - but is not equal to - a Gaussian density in a neighbourhood of the roots $\theta_{1,\star}$ and $\theta_{2,\star}$. Consequently, in such cases the limiting posterior shape is a Gaussian mixture with means given by $\theta_{1,\star}$ and $\theta_{2,\star}$, and variances $\Delta(\theta_{1,\star})^{-1}$ and $\Delta(\theta_{2,\star})^{-1}$.
5 Similar Approaches and Alternatives
5.1 Similarities to Other Gibbs Posteriors
The Q-posterior resembles the Bayesian synthetic likelihood posterior suggested in Wood, (2010), and developed in Price et al., (2018) (see, also, Frazier et al., , 2022). In particular, since $\theta_{n}=\arg_{\theta\in\Theta}\{0=\overline{m}_{n}(\theta)\}$, it follows that $\pi(\theta\mid\mathcal{Q}_{n})$ is equivalently represented as
$\pi(\theta\mid\mathcal{Q}_{n})=\frac{N\{\overline{m}_{n}(\theta_{n});0,W_{n}(\theta)/n\}\pi(\theta)}{\int_{\Theta}N\{\overline{m}_{n}(\theta_{n});0,W_{n}(\theta)/n\}\pi(\theta)\mathrm{d}\theta}.$
The above allows us to interpret the Q-posterior as a type of synthetic likelihood posterior based on observable summary statistic $\overline{m}_{n}(\theta_{n})$, which satisfies $0=\overline{m}_{n}(\theta_{n})$ by design, but where instead of simulating the mean and variance of the statistic we use the known mean and variance functions $\overline{m}_{n}(\theta)$ and $W_{n}(\theta)/n$.
The above interpretation of the Q-posterior also leads to efficient computation if we can calculate $\overline{m}_{n}(\theta)$ and $W_{n}(\theta)$ for any value of $\theta$ since samples from $\pi(\theta\mid\mathcal{Q}_{n})$ can then be obtained using standard MCMC tools. Further, in most cases, $W_{n}(\theta)$ can be taken as the sample variance of $\overline{m}_{n}(\theta)$, which can be calculated once $\overline{m}_{n}(\theta)$ is obtained (either directly or via automatic differentiation). In all numerical experiments herein we use random-walk Metropolis Hastings MCMC to produce samples from $\pi(\theta\mid\mathcal{Q}_{n})$ but note that stochastic gradient-based versions can also be used when it is cheap to compute $\nabla_{\theta}Q_{n}(\theta)$.
Chernozhukov and Hong, (2003) propose a type of generalized Bayesian inference based on either a set of over-identified estimation equations for $\theta$, i.e., more equations than unknown parameters, by taking a quadratic form of a vector of sample moments (for a related approach see Chib et al., , 2018), or by replacing the likelihood altogether with an M-estimator criterion; the latter is also used in a decision theoretic framework by Bissiri et al., (2016) to produce their generalized posterior. Philosophically, the approaches of Chernozhukov and Hong, (2003) and Bissiri et al., (2016) are based on conducting a form of Bayesian inference that bypasses the likelihood. In contrast, our approach is entirely motivated by attempting to produce posteriors with appropriate uncertainty quantification (see Section 5.2 for further discussion). Indeed, Chernozhukov and Hong, (2003) acknowledge in the discussion after their Theorem 2 that posteriors based on an M-estimator-like criterion function will not deliver calibrated posteriors in general, and the authors do not suggest how to rectify this issue. We also note that unlike existing generalized Bayes methods the Q-posterior does not require the difficulty of tuning a learning rate, which can greatly impact the posterior uncertainty quantification produced using generalized Bayesian methods.
5.2 Alternatives
In a likelihood context, when the model is correctly specified $\Sigma^{-1}=W(\theta_{\star})^{-1}=\mathcal{I}^{-1}_{\star}$, and the Q-posterior and exact posterior agree asymptotically. If the model is misspecified, the Q-posterior still yields credible sets that are calibrated, so long as $W_{n}(\theta_{\star})$ is a consistent estimator of $\mathcal{I}_{\star}$. Section 4 suggests that in most cases reliable estimators of $W(\theta_{\star})$ are available using existing formulas, robust estimators, or bootstrapping methods.
The Q-posterior approach is a significant departure from existing approaches to Bayesian inference in misspecified models. Two approaches that have so far received meaningful attention are the ‘sandwich’ correction suggested in Müller, (2013), and the BayesBag approach of Huggins and Miller, (2019).
Müller, ’s approach amounts to correcting the draws from the standard posterior using an explicit Gaussian approximation $\theta\sim N\{\bar{\theta},\mathcal{H}_{n}(\bar{\theta})^{-1}W_{n}(\bar{\theta})\mathcal{H}_{n}(\bar{\theta})^{-1}\}$, where $\bar{\theta}$ is the posterior mean. Such a correction can be implemented either by drawing directly from a multivariate normal, or by taking each posterior draw $\theta$ and modifying it according to the linear equation
$\widetilde{\theta}=\bar{\theta}+\mathcal{H}_{n}(\bar{\theta})^{-1}W_{n}(\bar{\theta})\mathcal{H}_{n}(\bar{\theta})^{1/2}(\theta-\bar{\theta});$
see Giummolè et al., (2019) for a related approach in the case of generalized posteriors built using scoring rules. We argue that this ex-post correction is sub-optimal for several reasons: first, philosophically, it amounts to the subsequent application of additional inference methods to the output of a Bayesian learning algorithm, and has no representation as a belief update in the sense of (1); second, it requires the explicit calculation of second-derivatives, which can be difficult and may be ill-behaved; third, this Gaussian approximation is poor when posteriors are not Gaussian, e.g., when the parameters have restricted support, in small sample sizes, or when the posterior is multi-modal; fourth, without additional constraints, this correction can easily produce a value of $\tilde{\theta}$ lying outside the support of $\pi(\theta)$, for instance, when $\Theta$ is a bounded subset of $\mathbb{R}^{d_{\theta}}$.
Posterior bagging, as suggested in the BayesBag approach (Huggins and Miller, , 2019) is an alternative method that attempts to correct posterior coverage through bagging. Letting $b=1,\dots,B$ denote bootstrap indices, and $y^{(b)}=(y_{1}^{(b)},\dots,y_{n}^{(b)})$ the $b$-th bootstrap sample, where $y_{i}^{(b)}$ is sampled with replacement from the original dataset. The BayesBag posterior is given by
$\pi^{\star}(\theta\mid y)\approx B^{-1}\sum_{b=1}^{B}\pi(\theta\mid y^{(b)}).$
BayesBag is easy to use, but requires re-running the MCMC sampling algorithm to obtain posterior draws of $\theta$ for each $\{y^{(b)}:b=1,\dots,B\}$.
In comparison to the Q-posterior, the BayesBag posterior has the following drawbacks. First, Huggins and Miller, (2019) demonstrate that the credible sets of the BayesBag posterior are not calibrated in general. Hence, the BayesBag posterior does not have valid frequentist coverage in general. Second, re-running the MCMC sampling algorithm to obtain posterior draws of $\theta$ for each $\{y^{(b)}:b=1,\dots,B\}$ can be computationally intensive. Third, applying the BayesBag approach to weakly dependent data, which are easily handled by the Q-posterior (see Section A.2.2 for one such example), is not straightforward.
Lyddon et al., (2019) propose an alternative posterior bootstrapping approach that attempts to correct the coverage of the posterior. Their approach is based on attempting to choose the learning rate $\omega$ in the generalized posterior to equate the information in this posterior with a loss-likelihood bootstrapped version, which has correct asymptotic coverage. While the resulting choice of $\omega$ equates the Fisher information number between the generalized posterior and its loss-bootstrapped version, it does not in general deliver calibrated posteriors.
6 Conclusion
We propose a new approach to Bayesian inference, which we call the Q-posterior, that delivers reliable uncertainty quantification. In likelihood-based settings the Q-posterior can be thought of as a type of Bayesian synthetic likelihood (Wood, , 2010) posterior where we replace the likelihood for the observed sample with an approximation for the likelihood of the score equations. The critical feature of the Q-posterior is that it delivers calibrated Bayesian inference regardless of model specification, while if the model is correctly specified the Q-posterior agrees with the exact posterior (in large samples).
When applied to generalized Bayesian posteriors (Bissiri et al., , 2016), the Q-posterior remains calibrated. All existing approaches of which we are aware attempt to correct the coverage of generalized posteriors using either ex-post correction of the posterior draws, which are ultimately based on some (implicit) normality assumption on the resulting posterior draws, or bootstrapping approaches. In contrast, the Q-posterior delivers correct uncertainty quantification without needing any additional tuning or ex-post correction of the draws.
When the likelihood is intractable and must be estimated, a version of the Q-posterior can still be implemented. However, this work uses importance sampling, or sequential importance sampling, along with Fisher’s identity to estimate the gradients used in the Q-posterior, which can then be used in MCMC schemes to produce posterior inference. In such cases, analyzing the behavior of the posterior becomes more difficult than the case where the scores are not estimated, and obtaining theoretical results similar to those in Theorem 1 is more onerous. Given the additional technicalities that are required to implement such an approach, we leave this topic for future research.
The empirical findings obtained by applying the Q-posterior to Tukey’s loss (Section A.2.3) suggest that the Q-posterior would also deliver reliable uncertainty quantification in empirical Bayes (see Casella, , 1985 for an introduction) settings where preliminary hyper-parameters are estimated before inference is conducted. We leave a detailed study of this conjecture for future research.
Acknowledgments
David Frazier was supported by the Australian Research Council’s Discovery Early Career Researcher Award funding scheme (DE200101070). Christopher Drovandi was supported by the Australian Research Council Future Fellowships Scheme (FT210100260). The authors thank Jeremias Knoblauch, and Jonathan Huggins for comments on a previous draft. All remaining errors are our own.
Appendix A Supplementary Material
A.1 Additional Numerical Results
A.1.1 Normal Linear Regression
This section contains complete numerical results for the case of $d=20$ covariates for the normal linear regression model considered in Section 3.1 of the main text. Please refer to Section 3.1 for full details on the numerical experiments.
Q-posterior Exact Bayes HrB
Bias Var Cover Bias Var Cover Bias Var Cover
$\beta_{1}$0.0003 0.001 0.97 -0.0002 0.001 0.98 0.0003 0.0017 0.99
$\beta_{2}$0.0012 0.001 0.97 -0.0009 0.001 0.96 -0.0023 0.0017 0.98
$\beta_{3}$0.002 0.001 0.95 -0.0019 0.001 0.95 -0.0019 0.0017 0.96
$\beta_{4}$ 0.0005 0.0011 0.95 0.0004 0.001 0.94 0.0012 0.0017 0.97
$\beta_{5}$0.0043 0.001 0.96 -0.0044 0.001 0.97 -0.0041 0.0018 0.98
$\beta_{6}$0.001 0.001 0.92 -0.0008 0.001 0.93 -0.0009 0.0017 0.96
$\beta_{7}$0.0006 0.001 0.94 -0.0004 0.001 0.94 -0.0005 0.0019 0.97
$\beta_{8}$0.0004 0.001 0.97 -0.0002 0.001 0.97 0.0005 0.0017 0.97
$\beta_{9}$0.0009 0.001 0.95 -0.0009 0.001 0.95 -0.0023 0.0018 0.98
$\beta_{10}$0.005 0.001 0.92 -0.005 0.001 0.93 -0.0059 0.002 0.95
$\beta_{11}$0.004 0.001 0.96 -0.004 0.001 0.96 -0.0031 0.0017 0.98
$\beta_{12}$ 0.0087 0.001 0.91 0.0089 0.001 0.91 0.0079 0.0017 0.92
$\beta_{13}$0.0011 0.001 0.98 -0.0012 0.001 0.96 -0.0005 0.0017 0.99
$\beta_{14}$ 0.0012 0.001 0.98 0.0016 0.001 0.98 0.0012 0.0018 0.98
$\beta_{15}$0.0094 0.001 0.99 -0.0093 0.001 0.99 -0.0098 0.0018 0.97
$\beta_{16}$ 0.0012 0.0011 0.95 0.0012 0.001 0.94 0.0025 0.0018 0.95
$\beta_{17}$0.0026 0.001 0.97 -0.0025 0.001 0.97 -0.0028 0.0018 0.96
$\beta_{18}$0.0032 0.001 0.94 -0.0034 0.001 0.93 -0.0024 0.0018 0.95
$\beta_{19}$0.0008 0.001 0.98 -0.0014 0.001 0.97 -0.003 0.0018 0.98
$\beta_{20}$ 0.0035 0.001 0.95 0.0035 0.001 0.95 0.0037 0.0018 0.99
Table 6: Accuracy results in the normal linear regression model for exact Bayes, HrB-posterior and Q-posterior when $d=20$ and $\gamma=0$ (i.e. correctly specified). Bias is the bias of the posterior mean across the replications. Var is the average posterior variance deviation across the replications. Cover is the actual posterior coverage, where the nominal level is set to 95% for the experiments.
Q-posterior Exact Bayes HrB
Bias Var Cover Bias Var Cover Bias Var Cover
$\beta_{1}$0.0004 0.001 0.97 -0.0004 0.001 0.97 -0.001 0.0015 0.97
$\beta_{2}$0.0024 0.0017 0.95 -0.002 0.001 0.87 0.0002 0.002 0.96
$\beta_{3}$0.0029 0.0017 0.95 -0.0026 0.001 0.85 -0.0006 0.0019 0.92
$\beta_{4}$ 0.0013 0.0011 0.97 0.0012 0.001 0.95 -0.0008 0.0017 0.98
$\beta_{5}$0.0043 0.0011 0.97 -0.0044 0.001 0.97 -0.0034 0.0017 0.97
$\beta_{6}$0.0012 0.001 0.96 -0.0011 0.001 0.95 0.0007 0.0018 0.95
$\beta_{7}$0.0008 0.001 0.95 -0.0006 0.001 0.96 -0.0009 0.0018 0.99
$\beta_{8}$0.0005 0.001 0.95 -0.0002 0.001 0.93 -0.0019 0.0016 0.99
$\beta_{9}$0.0003 0.001 0.95 -0.0002 0.001 0.96 0.0001 0.0017 0.96
$\beta_{10}$0.005 0.001 0.95 -0.005 0.001 0.94 -0.0055 0.0016 0.97
$\beta_{11}$0.0042 0.001 0.95 -0.0043 0.001 0.95 -0.0045 0.0017 0.99
$\beta_{12}$ 0.0088 0.0011 0.92 0.0088 0.001 0.91 0.0062 0.0017 0.96
$\beta_{13}$0.0006 0.001 0.98 -0.0006 0.001 0.97 -0.0025 0.0017 1.00
$\beta_{14}$ 0.0008 0.001 0.98 0.0011 0.001 0.98 0.0018 0.0016 1.00
$\beta_{15}$0.0087 0.001 0.97 -0.0086 0.001 0.96 -0.0102 0.0017 0.99
$\beta_{16}$ 0.0022 0.0011 0.92 0.0021 0.001 0.92 -0.0012 0.0016 0.98
$\beta_{17}$0.0009 0.001 0.98 -0.0008 0.001 0.98 -0.0019 0.0017 0.99
$\beta_{18}$0.0033 0.001 0.95 -0.0037 0.001 0.92 -0.0032 0.0016 0.96
$\beta_{19}$ 0.0005 0.001 0.99 0.0005 0.001 0.99 -0.0005 0.0016 0.98
$\beta_{20}$ 0.0044 0.001 0.98 0.0043 0.001 0.97 0.0039 0.0018 0.97
Table 7: Accuracy results in the normal linear regression model for exact Bayes, HrB-posterior and Q-posterior when $d=20$ and $\gamma=2$ (i.e. misspecified). Bias is the bias of the posterior mean across the replications. Var is the average posterior variance deviation across the replications. Cover is the actual posterior coverage, where the nominal level is set to 95% for the experiments.
A.1.2 Additional Results Poisson Regression
This section contains complete numerical results for the case of $d_{\theta}=10$ and $d_{\theta}=20$ covariates for the Poisson regression model considered in Section 3.2 of the main text. Please refer to Section 3.2 for full details on the numerical experiments.
Q-posterior Exact Bayes GenBayes
Bias Var Cover Bias Var Cover Bias Var Cover
$\theta_{1}$1.44 0.055 0.92 -0.915 0.037 0.88 -0.945 0.055 0.93
$\theta_{2}$ 1.36 0.032 0.96 1.319 0.022 0.94 1.317 0.032 0.97
$\theta_{3}$0.45 0.032 0.95 -0.407 0.021 0.88 -0.398 0.032 0.95
$\theta_{4}$0.16 0.032 0.95 -0.219 0.022 0.88 -0.222 0.032 0.98
$\theta_{5}$0.40 0.031 0.93 -0.392 0.021 0.89 -0.392 0.032 0.95
$\theta_{6}$ 0.59 0.032 0.96 0.594 0.021 0.92 0.596 0.032 0.96
$\theta_{7}$ 0.08 0.031 0.94 0.087 0.021 0.91 0.080 0.032 0.94
$\theta_{8}$ 0.74 0.031 0.97 0.747 0.021 0.93 0.752 0.031 0.96
$\theta_{9}$ 0.26 0.032 0.98 0.273 0.022 0.95 0.275 0.032 0.97
$\theta_{10}$ 0.03 0.032 0.95 0.037 0.021 0.86 0.028 0.032 0.94
Table 8: Accuracy results in the Poisson regression model for exact Bayes, Q-posterior, and generalized Bayes (GenBayes) based on the quasi-likelihood. Bias is the bias of the posterior mean across the replications and has been multiplied by 1000 for readability. Var is the average posterior variance deviation across the replications and has been multiplied by 100 for readability. Cover is the actual posterior coverage, where the nominal level is set to 95% for the experiments.
Q-posterior Exact Bayes GenBayes
Bias Var Cover Bias Var Cover Bias Var Cover
$\theta_{1}$1.769 0.056 0.95 -1.024 0.037 0.90 -1.125 0.056 0.97
$\theta_{2}$0.469 0.033 0.94 -0.476 0.022 0.88 -0.481 0.034 0.95
$\theta_{3}$ 0.245 0.033 0.96 0.254 0.022 0.93 0.251 0.033 0.96
$\theta_{4}$ 0.550 0.033 0.96 0.547 0.022 0.92 0.534 0.034 0.96
$\theta_{5}$ 0.662 0.033 0.94 0.663 0.022 0.89 0.665 0.033 0.94
$\theta_{6}$ 0.260 0.033 1.00 0.276 0.022 0.94 0.278 0.033 0.99
$\theta_{7}$ 0.644 0.033 0.98 0.630 0.022 0.93 0.628 0.033 0.98
$\theta_{8}$0.334 0.032 0.97 -0.333 0.022 0.89 -0.362 0.033 0.97
$\theta_{9}$ 0.593 0.032 0.94 0.615 0.022 0.91 0.609 0.033 0.94
$\theta_{10}$0.601 0.033 0.96 -0.604 0.022 0.92 -0.589 0.033 0.95
$\theta_{11}$0.677 0.033 0.98 -0.671 0.022 0.95 -0.678 0.033 0.97
$\theta_{12}$0.079 0.033 0.98 -0.076 0.022 0.97 -0.084 0.033 0.98
$\theta_{13}$ 0.409 0.033 0.99 0.430 0.022 0.95 0.418 0.033 0.98
$\theta_{14}$ 1.275 0.033 0.98 1.260 0.022 0.95 1.281 0.033 0.98
$\theta_{15}$ 0.065 0.033 0.93 0.073 0.022 0.88 0.077 0.033 0.92
$\theta_{16}$0.229 0.032 0.95 -0.219 0.022 0.90 -0.216 0.033 0.94
$\theta_{17}$ 0.460 0.033 0.95 0.458 0.022 0.92 0.474 0.033 0.95
$\theta_{18}$0.795 0.033 0.95 -0.793 0.022 0.89 -0.780 0.033 0.94
$\theta_{19}$0.036 0.033 0.96 -0.023 0.022 0.95 -0.023 0.033 0.96
$\theta_{20}$ 0.448 0.033 0.98 0.440 0.022 0.94 0.445 0.033 0.98
Table 9: Accuracy results in the Poisson regression model for exact Bayes, Q-posterior, and generalized Bayes (GenBayes) based on the quasi-likelihood. Bias is the bias of the posterior mean across the replications and has been multiplied by 1000 for readability. Var is the average posterior variance deviation across the replications and has been multiplied by 100 for readability. Cover is the actual posterior coverage, where the nominal level is set to 95% for the experiments.
A.2 Additional Examples
A.2.1 Robust Location Inference
Consider observing a sequence $y_{1},\dots,y_{n}$ from the model
$y_{i}=\theta+\epsilon_{i},\quad i=1,\dots,u,\quad\epsilon_{i}\stackrel{{\scriptstyle iid}}{{\sim}}F(\cdot)$
where $F(\cdot)$ is unknown, the unknown parameter $\theta$ has prior density $\pi(\theta)$ and is independent of $\epsilon$. Our goal is posterior inference on $\theta$, and for reasons of robustness we follow Doksum and Lo, (1990), and consider posterior inference based on the sample median $T_{n}=\text{med}(y_{1},\dots,y_{n})$, which is known to be robust to outliers in the data; for a related approach see Lewis et al., (2021). When $F$ has density $f$, and $n$ is odd, Doksum and Lo, (1990) show that the posterior $\pi(\theta\mid T_{n})$ is
$\pi(\theta\mid T_{n})=\pi(\theta)\exp\left\{\frac{1}{2}(n-1)\log F(T_{n}-\theta)\{1-F(T_{n}-\theta)\}+\log f_{\theta}(T_{n})(1-F(T_{n}-\theta)))\right\}.$
Given the above form of the posterior, Miller, (2021) suggests using generalized Bayesian inference for $\theta$ with the simpler loss function $\mathcal{D}_{n}(\theta)=-\frac{1}{2}\log F(T_{n}-\theta)\{1-F(T_{n}-\theta)\}$. While Miller, (2021) demonstrates that such a posterior is well-behaved in large samples, the resulting posterior does not have correct coverage even when the model is correctly specified. It is simple to show that the same result, i.e., inaccurate posterior coverage, applies if one uses the original posterior $\pi(\theta\mid T_{n})$ proposed in Doksum and Lo, (1990), or the more general approach suggested in Lewis et al., (2021).
In contrast, if we conduct inference using the Q-posterior based on the loss $\mathcal{D}_{n}(\theta)=\ell_{n}(\theta)$, then the resulting posterior is well-behaved and correctly quantifies uncertainty. The Q-posterior requires calculating the gradient of $\mathcal{D}_{n}(\theta)$ with respect to $\theta$, given by
$\overline{m}_{n}(\theta)=\frac{1}{2n}\frac{f(T_{n}-\theta)}{F(T_{n}-\theta)}-\frac{1}{2}\frac{f(T_{n}-\theta)}{1-F(T_{n}-\theta)},$
and the variance of $\sqrt{n}\overline{m}_{n}(\theta)$, for any $\theta\in\Theta$. For $W_{n}(\theta)$ denoting an estimator of $\text{var}\{\sqrt{n}\overline{m}_{n}(\theta)\}$, the Q-posterior is
$\pi(\theta\mid\mathcal{Q}_{n})\propto\pi(\theta)\exp\left[-\frac{1}{2n}\left\{\frac{1}{2}\frac{f(T_{n}-\theta)}{F(T_{n}-\theta)}-\frac{1}{2}\frac{f(T_{n}-\theta)}{1-F(T_{n}-\theta)}\right\}^{2}/W_{n}(\theta)\right].$
A simple estimator for $\text{var}\{\sqrt{n}\overline{m}_{n}(\theta)\}$ can be obtained using bootstrap replications of the median, denoted as $\{T_{n}^{(b)}:b=1,\cdots,B\}$. Such an estimator is extremely fast since $\{T_{n}^{(b)}:b=1,\cdots,B\}$ are only generated once outside of the sampling algorithm. For any value of $\theta$, we can then take $W_{n}(\theta)$ as the sample variance of the observations $\{\overline{m}_{n}(\theta){|}_{T_{n}=T_{n}^{(b)}}:b=1,\dots,B\}$, which only requires evaluating the closed-form gradient $B$ times and then taking the sample variance of the evaluations. [^5] Such an approach is similar to the estimating equations bootstrap, see Hu and Kalbfleisch, , 2000 and Chatterjee and Bose, , 2005 for a discussion. Consequently, the steps necessary to calculate $W_{n}(\theta)$ are precisely the same as forming any sample variance estimator.
We now compare the uncertainty quantification produced using the generalized and Q-posteriors in correctly and misspecified models. In both cases, we assume $F(\cdot)$ is a standard Gaussian distribution. In the first experiment, referred to as DGP1, the observed data is generated from a Gaussian distribution with mean $\theta=1$, and variance 4. In the misspecified regime, referred to as DGP2, we generate observed data from a mixed Gaussian distribution with parameterization $0.9N(\theta=1,4)+0.1N(0,1).$
In the first case the true median is unity, while in the second case the true median of the data is approximately $0.84$; this value is found numerically by inverting the CDF of the mixture distribution. We simulate 1000 observed datasets from both DGPs and compare the results of generalized Bayes and that based on the Q-posterior. Table 10 compares the posterior means, variances and coverage of the generalized Bayes and the Q-posterior procedures. The results demonstrate that the generalized posterior proposed in Miller, (2021) does not produce reliable coverage for the true median, while the coverage of the Q-posterior is again close to the nominal level.
Q-posterior GenBayes
DGP1 Bias Var Cover Bias Var Cover
$\theta$0.0031 0.0654 0.9420 -0.0031 0.0152 0.6700
DGP2 Bias Var Cover Bias Var Cover
$\theta$0.0115 0.0607 0.9360 -0.0118 0.0157 0.7160
Table 10: Posterior accuracy results for the median using the generalized posterior (GenBayes) and Q-posterior. Bias is the bias of the posterior mean across the replications. Var is the average posterior variance deviation across the replications. Cover is the posterior coverage (95% nominal coverage). For GenBayes we set $\omega=1$ in all experiments.
A.2.2 Approximate Inference in Time Series Models
This example shows how the Q-posterior can provide calibrated Bayesian inferences when the assumed data generating process is correct, but an approximate likelihood is used to speed-up computation. Furthermore, this example highlights how automatic differentiation can be used to calculate the scores that are not easily accessible analytically.
Let $\{X_{t}\}_{t=1}^{n}$ be a discretely observed, zero-mean, random variable generated according to an autoregressive fractionally integrated moving average (ARFIMA$(2,d,1)$) model
$(1-\phi_{1}L-\phi_{2}L^{2})^{d}X_{t}=(1-\vartheta_{1}L)\epsilon_{t},\quad\theta=(\phi_{1},\phi_{2},\vartheta_{1},d)^{\top},$
where $L$ is the lag operator, and $\epsilon_{t}\sim\mathcal{N}(0,\sigma^{2})$. We consider observed data generated from the ARFIMA($2,d,1$) model with true parameter $\theta_{\star}=(\phi_{1\star},\phi_{2\star},\vartheta_{1,\star},d_{\star})^{\top}=(0.45,0.1,-0.4,0.4)^{\top}$ and we set $n=15,000$.
For simplicity, we impose independent uniform priors on the components of $\theta$, i.e., for $\mathcal{U}(a,b)$ denoting a uniform random variable with support $(a,b)$, $\phi_{1}\sim\mathcal{U}(-1,1)$, $\phi_{2}\sim\mathcal{U}(-1,1)$, $\vartheta_{1}\sim\mathcal{U}(-1,1)$ and $d\sim\mathcal{U}(-0.5,0.5)$. Bon et al., (2021) impose stationarity constraints on the parameters, but given the large sample size investigated here, the prior has minimal influence.
The likelihood function of the ARFIMA model for large $n$ is computationally intensive (Chan and Palma, , 1998). Consequently, we approximate the likelihood by the Whittle likelihood (Whittle, , 1953) to form an approximate posterior; see, also, Salomone et al., (2020) and Bon et al., (2021) for a similar approach. We compare the approximate posterior and the Q-posterior based on the Whittle likelihood, and show that the Q-posterior delivers reliable uncertainty quantification, while the approximate posterior has poor uncertainty quantification.
The Whittle likelihood operates with the data and covariance function in the frequency domain, as opposed to the time domain, and is based on the spectral density of the model, and the periodogram. For a frequency $\omega\in[-\pi,\pi]$, the spectral density of the ARFIMA($2,d,1$) for $d\in(-0.5,0.5)$ is
$f_{\theta}(\omega)=\frac{\sigma^{2}}{2\pi}\frac{|1-\vartheta_{1}\exp(i\omega)|^{2}}{\left|\left\{1-\phi_{1}\exp(i\omega)-\phi_{2}\exp(2i\omega)\right\}\{1-\exp(i\omega)\}^{d}\right|^{2}}.$
The periodigram of the data at the frequencies $\{\omega_{k}:k=-\lceil n/2\rceil+1,\ldots,\lfloor n/2\rfloor\}$ is
$\mathcal{F}(\omega_{k})=\frac{|J(\omega_{k})|^{2}}{n},\quad J(\omega_{k})=\frac{1}{\sqrt{2\pi}}\sum_{t=1}^{n}X_{t}\exp(-i\omega_{k}t),\quad\omega_{k}=\frac{2\pi k}{n},$
where $J(\omega_{k})$ is the discrete Fourier transform (DFT) of the time series. The Whittle log-likelihood for zero-mean data is then (Whittle, , 1953)
$\ell_{\text{whittle}}(\theta)=-\sum_{k=-\lceil n/2\rceil+1}^{\lfloor n/2\rfloor}\left(\log f_{\theta}(\omega_{k})+\frac{\mathcal{F}(\omega_{k})}{f_{\theta}(\omega_{k})}\right).$
In practice, it is only necessary to calculate the summation over about half of the Fourier frequencies, $\omega_{k}$, due to symmetry about $\omega_{0}=0$ and since $f_{\theta}(\omega_{0})=0$ for centred data.
The periodogram can be calculated in $\mathcal{O}(n\log n)$ time, and only needs to be calculated once per dataset. Given the periodogram, the cost of each subsequent likelihood evaluation is $\mathcal{O}(n)$, compared to the usual likelihood cost for time series (with dense precision matrix) which is $\mathcal{O}(n^{2})$.
The Whittle likelihood effectively treats $\mathcal{F}(\omega_{k})$ for each $\omega_{k}$ as being an independent exponential random variate with mean $f_{\theta}(\omega)$. Given the assumed independence, it makes the Q-posterior easy to apply in principle. However, the score for each component of $\ell_{\text{whittle}}(\theta)$ is not easy to obtain analytically. Herein, we use automatic differentiation in Julia (Bezanson et al., , 2017), specifically the ForwardDiff.jl library (Revels et al., , 2016) to compute the scores $\overline{m}_{n}(\theta)$.
We use the AdaptiveMCMC.jl library (Vihola, , 2014) to draw samples from the posterior under the Whittle approximation and the Q-posterior approach. So long as the model is correctly specified, and under appropriate regularity conditions and restrictions on the true parameters, point-estimates obtained using the Whittle likelihood converge asymptotically to the true value generating the data (see, e.g., Fox and Taqqu, , 1986, and Dahlhaus, , 1989). Thus, given the large sample size considered here, the pseudo-true parameter and the true parameter coincide. Table 11 gives accuracy measures based on 100 independent datasets (we use the arfima package (Veenstra, , 2012) in R (R Core Team, , 2022) to simulate datasets). The results show that the Whittle approximation has low bias and MSE, but produces posterior approximations that are over-concentrated, as the poor coverage rates show. In contrast, the Q-posterior approximations exhibit substantially more accurate uncertainty quantification, but still having biases that are similar to those based on the Whittle likelihood.
Method MSE Bias St. Dev. Coverage (90%)
$\phi_{1}$
Whittle 0.0028 -0.0055 0.030 72
Q-posterior 0.0038 -0.0024 0.044 93
$\phi_{2}$
Whittle 0.0013 0.00031 0.021 74
Q-posterior 0.0017 -0.00038 0.029 91
$\vartheta_{1}$
Whittle 0.0021 -0.0011 0.026 73
Q-posterior 0.0027 0.00024 0.037 89
$d$
Whittle 0.00078 0.0027 0.016 72
Q-posterior 0.0012 0.00061 0.025 88
Table 11: Results for the Whittle example using 100 independent datasets simulated with true parameter value $\theta=(\phi_{1},\phi_{2},\vartheta_{1},d)^{\top}=(0.45,0.1,-0.4,0.4)^{\top}$. MSE denotes mean squared error, Bias the average bias, St.Dev. the average standard deviation, and Coverage is the actual coverage rate of the credible sets.
In Figure 3 we compare the approximate posteriors for the Whittle likelihood and Q-posterior for a single dataset, the same one used in Bon et al., (2022) generated with the same true parameter considered here. We also compare the approximate posteriors with the true posterior generated in Bon et al., (2022). It is evident that the posterior approximations based on the Whittle likelihood are overconcentrated compared to the true posterior. In contrast, the Q-posteriors are quite close to the true posteriors for this particular dataset.
Figure 3: Univariate densities estimates of approximations to the posterior distribution for a single dataset generated from an ARFIMA with true parameter value $\theta=(\phi_{1},\phi_{2},\vartheta_{1},d)^{\top}=(0.45,0.1,-0.4,0.4)^{\top}$. Shown are posterior approximations based on the Whittle likelihood (orange), Q-posterior (green) and the exact likelihood (blue).
On a laptop computer with an Intel i7-12800H processor (3.70 GHz clock speed) and 32GB RAM the approximate timing for a single likelihood evaluation with the Whittle likelihood, Q-score and true likelihood are 0.001, 0.01 and 0.17 seconds. The true likelihood uses the arfima package in R.
A.2.3 Tukey’s Loss: Contaminated normal example
In certain cases our preferred loss function depends on both unknown parameters, $\theta$, and unknown hyper-parameters, $\kappa$. In such settings, Jewson and Rossell, (2022) propose using the Hyvärinen score for Bayesian inference and model selection on the unknown hyperparameter $\kappa$. They essentially propose a two-step procedure one first learns, or choose, the hyper-parameters $\kappa$ using a Gibbs posterior based on the Hyvärinen score, and then conducts generalized Bayesian inference on the remaining parameter $\theta$ given the chosen value of the hyper-parameter. While such an approach is sensible, since inference on $\theta$ is based on a generalized posterior, the posterior that results in the second-step does not reliably quantify uncertainty.
In this example, the Q-posterior delivers reliable uncertainty quantification even when the loss-function in question depends on hyper-parameters that must be fixed/estimated before the loss can be computed. We show that even though these hyper-parameters are fixed in a first step, the Q-posterior still delivers reliable uncertainty quantification.
One of the motivating examples in Jewson and Rossell, (2022) is robust regression, where Tukey’s loss is considered. In the case of no covariates, Tukey’s loss for a single observation $y$ is given by
$q\left(y;\theta,\kappa\right)$ =$=\begin{cases}\frac{1}{2}\log\left(2\pi\phi\right)+\frac{\left(y-\mu\right)^{2}}{2\phi}-\frac{\left(y-\mu\right)^{4}}{2\phi^{2}\kappa^{2}}+\frac{\left(y-\mu\right)^{6}}{6\phi^{3}\kappa^{4}},&\text{ if }\left|y-\mu\right|\leq\kappa\sqrt{\phi}\\ \frac{1}{2}\log\left(2\pi\phi\right)+\frac{\kappa^{2}}{6},&\text{ otherwise }\end{cases}$
Tukey’s loss depends on the model parameters $\mu,\phi$, and the hyper-parameter $\kappa$, which controls the tails of the loss. The value of $\kappa$ artificially deflates extreme data points, such as gross outliers, to ensure that these points do not unduly influence inferences. The use of Tukey’s loss is complicated by the fact that if all three parameters, $\mu,\phi,\kappa$, are permitted to vary, the loss is minimized by driving $\kappa$ to 0, which is not meaningful.
We consider the same example as in Section 5.1 of Jewson and Rossell, (2022), where the true data generating process involves taking $n$ independent draws from a mixture model, $0.9\mathcal{N}(0,1)+0.1\mathcal{N}(5,3)$. The assumed model for the data is $\mathcal{N}(\mu,\phi)$, so that $\theta=(\mu,\phi)^{\top}$, and the assumed model is correct up to a small proportion of outliers. As in Jewson and Rossell, (2022), we consider datasets of size $n=500$, and set the priors as $\mu\sim\mathcal{N}(0,5)$ and $\phi\sim\mathrm{Exp}(1)$.
Section 3.4 of Jewson and Rossell, (2022) uses a two-step approach, whereby they first choose the value of $\kappa$ by maximizing the $\mathcal{H}$-score over $\kappa$, and then conduct generalized Bayes inference for the remaining parameters with a fixed value of $\kappa$ from the optimization. We consider a similar approach whereby we fix $\kappa$ and conduct generalized Bayes inference for $\mu$ and $\phi$ using Tukey’s loss. For computational simplicity, we set $\kappa=6$ for all experiments, since this value is close to the optimal value found in Jewson and Rossell, (2022), and then use the Q-posterior to conduct inference on $\mu$ and $\phi$.
Tukey’s loss has a continuous derivative in $\phi$ and $\mu$ except for sets of measure zero (i.e., it is equicontinuous), and can be shown to satisfy the maintained regularity conditions we later use to deduce the asymptotic behavior of the Q-posterior. Thus, we proceed by computing the score for each component of the loss, and return the relevant score at each iteration of the MCMC, depending on the condition stated in the loss $\left|x-\mu\right|\leq\kappa\sqrt{\phi}$; e.g., if this condition is satisfied we use the score associated to the first component in the loss, else we use the score of the second component.
By minimizing Tukey’s loss on a dataset of size 10 million with fixed hyper-parameter of $\kappa=6$, we find that the pseudo-true parameter value is given by $\theta^{*}\approx(0.088,1.00)^{\top}$, and we investigate if the Q-posterior posterior produces well-calibrated inference for this value of $\theta_{\star}$. Again we use the AdaptiveMCMC.jl library for sampling the Q-posterior posterior.
We compare our approach with generalized Bayesian inference based on Tukey’s loss, i.e., (1) based on Tukey’s loss, where we set the learning rate to be $\omega=1$ and $\omega=2$. Table 12 suggest that generalized Bayesian inference based on Tukey’s loss produces smaller mean squared error, but the Q-posterior yields much more accurate uncertainty quantification, especially for $\phi$. That is, since the squared biases are similar between the two methods, the difference in mean squared error is attributable to the smaller variance for the generalized posterior, which results in poor coverage. Figure 4 displays posterior approximations for a single dataset, and shows that the Q-posterior has additional variability relative to posteriors based on Tukey’s loss.
Method MSE Bias St. Dev. Coverage (90%)
$\mu$
GBI $\omega=1$ 0.0089 0.023 0.057 78
GBI $\omega=2$ 0.0073 0.022 0.040 65
Q-posterior 0.015 0.058 0.080 83
$\phi$
GBI $\omega=1$ 0.069 0.066 0.12 66
GBI $\omega=2$ 0.060 0.058 0.084 50
Q-posterior 0.18 0.21 0.27 86
Table 12: Results for the Q-posterior and generalized Bayesian inference (GBI) based on Tukey loss example using 100 independent datasets for $\kappa=6$. Here the pseudo-true parameter value is $\theta^{*}\approx(0.088,1.00)^{\top}$. For a detailed description on the column entries, see the caption in Table 11.
Figure 4: Univariate density estimates of approximations to the posterior distribution for a single dataset for the Tukey loss example with pseudo-true parameter value $\theta^{*}\approx(0.088,1.00)^{\top}$. Posterior approximations for the Q-posterior are shown in green and blue for generalised Bayesian inference (GBI).
A.3 Lemmas
[…]
The following lemmas are used to prove the main results. To simplify the statement and derivation of results, redefine $Q_{n}(\theta)$ as
$Q_{n}(\theta)=-\frac{n}{2}\overline{m}_{n}(\theta)^{\top}W_{n}(\theta)^{-1}\overline{m}_{n}(\theta).$
While this definition is not as easily interpretable as the $Q_{n}(\theta)$ defined in equation (2), it simplifies our manipulations.
Lemma 1.
Under Assumptions 1-5, the following are satisfied for some $\theta_{\star}\in\Theta_{\star}$, and $\theta_{n}:=\arg_{\theta\in\Theta}\{0=\overline{m}_{n}(\theta)\}$.
1. $\|\theta_{n}-\theta_{\star}\|=O_{p}(n^{-1/2})$.
2. For any $\delta_{n}=o(1)$, $T_{n}=\{\theta\in\Theta,\theta_{\star}\in\Theta_{\star}:\|\theta-\theta_{\star}\|\leq\delta_{n}\}$, and $t={\sqrt{n}}(\theta-\theta_{\star}),\;\theta\in T_{n}$,
$Q_{n}(\theta_{\star}+t/{\sqrt{n}})-Q_{n}(\theta_{\star})=t^{\top}\mathcal{H}_{n}(\theta_{\star})W_{n}(\theta_{\star})^{-1}\sqrt{n}\overline{m}_{n}(\theta_{\star})-\frac{1}{2}t^{\top}\Delta(\theta_{\star})t+R_{n}(\theta_{\star}+t/{\sqrt{n}}),$
for a remainder term $R_{n}(\theta)$ satisfying
$\sup_{\|\theta-\theta_{\star}\|\leq\delta_{n}}\frac{R_{n}(\theta)}{1+\|\sqrt{n}(\theta-\theta_{\star})\|^{2}}=o_{p}(1).$
Proof.
Proof of 1. From the consistency of $\theta_{n}$ there exists some positive $\delta_{n}=o(1)$ such that $P_{0}^{(n)}\left\{\|\theta_{n}-\theta_{\star}\|\geq\delta_{n}\right\}=o(1)$. By Assumption 1(ii), we have
$\sup_{\|\theta-\theta_{\star}\|\leq\delta_{n}}\|\overline{m}_{n}(\theta)-\overline{m}(\theta)-\overline{m}_{n}(\theta_{\star})\|=o_{p}(n^{-1/2})$ (5)
under $P^{(n)}_{0}$. Then, with $P^{(n)}_{0}$ probability converging to one for the sequence $\delta_{n}$, we have
$\|\overline{m}_{n}(\theta)-m(\theta_{n})-\overline{m}_{n}(\theta_{\star})\|$ $\leq o_{p}(n^{-1/2})$
$\|\overline{m}_{n}(\theta)-m(\theta_{n})-\overline{m}_{n}(\theta_{\star})\|$ $\geq\|m(\theta_{n})\|-\|\overline{m}_{n}(\theta)\|-\|\overline{m}_{n}(\theta_{\star})\|.$
Rearranging terms, and applying Assumption 1(iv),
$\|m(\theta_{n})\|$ $\leq o_{p}(n^{-1/2})+\|\overline{m}_{n}(\theta_{\star})\|\{1+o_{p}(1)\}=O_{p}(n^{-1/2}).$
Because $m(\theta)$ differentiable, $\mathcal{H}_{\star}=\mathcal{H}(\theta_{\star})$ is positive-definite - Assumption 1(iii) - there exists $C>0$ such that $C\|\theta_{n}-\theta_{\star}\|\leq\|m(\theta_{n})\|\leq O_{p}(n^{-1/2}).$
Proof of 2. On the set $T_{n}$, the result follows from the following expansion of $Q_{n}(\theta)=-\frac{n}{2}\overline{m}_{n}(\theta)^{\top}W_{n}(\theta)^{-1}\overline{m}_{n}(\theta)$ around $\theta_{\star}$:
$Q_{n}(\theta)-Q_{n}(\theta_{\star})={\sqrt{n}}(\theta-\theta_{\star})^{\prime}\mathcal{H}(\theta_{\star})^{\top}W(\theta_{\star})^{-1}{\sqrt{n}}\overline{m}_{n}(\theta_{\star})-{\sqrt{n}}(\theta-\theta_{\star})^{\prime}\Delta(\theta_{\star}){\sqrt{n}}(\theta-\theta_{\star})/2+R_{n}(\theta);$
where $R_{n}(\theta)=R_{1n}(\theta)+R_{3n}(\theta)+R_{3n}(\theta)$, and
$R_{1n}(\theta)=$ $-n\left\{\overline{m}_{n}\left(\theta_{\star}\right)^{\top}W(\theta)^{-1}\mathcal{H}\left(\theta_{\star}\right)\left(\theta-\theta_{\star}\right)-\frac{1}{2}\left(\theta-\theta_{\star}\right)^{\top}\mathcal{H}\left(\theta_{\star}\right)^{\prime}W(\theta)^{-1}\mathcal{H}\left(\theta_{\star}\right)\left(\theta-\theta_{\star}\right)\right\}$
+$+\left\{\frac{1}{2}\overline{m}_{n}\left(\theta\right)^{\top}W_{n}(\theta)^{-1}\overline{m}_{n}(\theta)+\frac{1}{2}\overline{m}_{n}\left(\theta_{\star}\right)^{\top}W_{n}(\theta)^{-1}m\left(\theta_{\star}\right)\right\}$
$R_{2n}(\theta)=$ $-n\left[\frac{1}{2}\overline{m}_{n}\left(\theta_{\star}\right)^{\top}\left\{W_{n}\left(\theta_{\star}\right)^{-1}-W_{n}(\theta)^{-1}\right\}\overline{m}_{n}\left(\theta_{\star}\right)\right]$
$R_{3n}(\theta)=$ $n\left[\overline{m}_{n}\left(\theta_{\star}\right)^{\top}\left\{W(\theta)^{-1}-W\left(\theta_{\star}\right)^{-1}\right\}\mathcal{H}\left(\theta_{\star}\right)\left(\theta-\theta_{\star}\right)\right]$
$-\left[\frac{1}{2}\left(\theta-\theta_{\star}\right)^{\prime}\mathcal{H}\left(\theta_{\star}\right)^{\prime}\left\{W(\theta)^{-1}-W\left(\theta_{\star}\right)^{-1}\right\}\mathcal{H}\left(\theta_{\star}\right)\left(\theta-\theta_{\star}\right)\right].$
The stated condition on $R_{n}(\theta)$ then directly follows by Lemma 3. ∎
Lemma 2.
Under Assumptions 1-5, with probability converging to one under $P^{(n)}_{0}$,
$0\leq{\pi(\theta_{n}\mid\mathcal{Q}_{n})}/{{n^{{d_{\theta}}/2}}}\leq{1}/\{{{(2\pi)^{d_{\theta}}|\Delta|}}\}^{1/2}.$
Proof.
The proof proceeds via a similar argument to that used to prove Lemma 1 in Frazier et al., (2021). Let $\gamma_{n}=o(1)$ with $\gamma_{n}{\sqrt{n}}\rightarrow\infty$, and let $\mathcal{N}_{\gamma_{n}}=\{\theta\in\Theta:\|\theta-\theta_{n}\|\leq\gamma_{n}\}$. Rewrite the posterior $\pi(\theta\mid\mathcal{Q}_{n})$ as
$\pi(\theta\mid\mathcal{Q}_{n})$ $=\pi(\theta_{n}\mid\mathcal{Q}_{n})\frac{M_{n}(\theta)\pi(\theta)}{M_{n}(\theta_{n})\pi(\theta_{n})}\exp\{Q_{n}(\theta)-Q_{n}(\theta_{n})\}$
$=\pi(\theta_{n}\mid\mathcal{Q}_{n})\exp\{Q_{n}(\theta)-Q_{n}(\theta_{n})\}+\pi(\theta_{n}\mid\mathcal{Q}_{n})\left\{\frac{M_{n}(\theta)\pi(\theta)}{M_{n}(\theta_{n})\pi(\theta_{n})}-1\right\}\exp\{Q_{n}(\theta)-Q_{n}(\theta_{n})\},$
where we remind the reader that $M_{n}(\theta)=|W_{n}(\theta)|^{-1/2}$. From Lemma 1(1), for any $\gamma_{n}=o(1)$ such that $\gamma_{n}\sqrt{n}\rightarrow+\infty$, and $\Omega_{n}:=\|\theta_{n}-\theta_{\star}\|\leq\gamma_{n}$ $P^{(n)}_{0}\left\{\Omega_{n}\right\}=1+o(1)$. On $\Omega_{n}$, by Assumptions 2 and 3, and the compactness of $\{\theta\in\Theta:\|\theta-\theta_{n}\|\leq\gamma_{n}\}$,
$\left\{\frac{M_{n}(\theta)\pi(\theta)}{M_{n}(\theta_{n})\pi(\theta_{n})}-1\right\}\leq\sup_{\|\theta-\theta_{n}\|\leq\gamma_{n}}\left|\left\{\frac{M_{n}(\theta)\pi(\theta)-M_{n}(\theta_{n})\pi(\theta_{n})}{M_{n}(\theta_{n})\pi(\theta_{n})}\right\}\right|=o_{p}(1).$
Hence, over $\mathcal{N}_{\gamma_{n}}$,
$\pi(\theta\mid\mathcal{Q}_{n})$ $=\pi(\theta_{n}\mid\mathcal{Q}_{n})\{1+o_{p}(1)\}\exp\{Q_{n}(\theta)-Q_{n}(\theta_{n})\},$
and
$\int_{\mathcal{N}_{\gamma_{n}}}\pi(\theta\mid\mathcal{Q}_{n})\mathrm{d}\theta$ $=\pi(\theta_{n}\mid\mathcal{Q}_{n})\{1+o_{p}(1)\}\int_{\mathcal{N}_{\gamma_{n}}}\exp\left\{Q_{n}(\theta)-Q_{n}(\theta_{n})\right\}\mathrm{d}\theta$
$=\frac{\pi(\theta_{n}\mid\mathcal{Q}_{n})}{{n^{d_{\theta}/2}}}\{1+o_{p}(1)\}\int_{\|t\|\leq\gamma_{n}{\sqrt{n}}}\exp\{Q_{n}(\theta_{n}+t/{\sqrt{n}})-Q_{n}(\theta_{n})\}\mathrm{d}t,$ (6)
where the second line follows from the change of variables $t=\sqrt{n}(\theta-\theta_{n})$.
From the expansion in Lemma 1, over the set $\mathcal{N}_{\gamma_{n}}$, for $\Delta(\theta)=\mathcal{H}(\theta)\mathcal{I}(\theta)^{-1}\mathcal{H}(\theta)$,
$Q_{n}(\theta_{n}+t/{\sqrt{n}})-Q_{n}(\theta_{n})$ $=-\frac{1}{2}t^{\top}\Delta(\theta_{n})t+O(\|t\|^{2}\gamma_{n})=-\frac{1}{2}t^{\top}\Delta(\theta_{n})t+o(1).$
Rewrite the above as
$Q_{n}(\theta_{n}+t/{\sqrt{n}})-Q_{n}(\theta_{n})$ $=-\frac{1}{2}t^{\top}\left(I+V_{n}\right)\Delta(\theta_{\star})t+o(1),$ (7)
where $V_{n}=\left[\Delta(\theta_{n})-\Delta(\theta_{\star})\right]\Delta(\theta_{\star})^{-1}$. By Assumptions 1(iii) and 2, for some $C>0$, $\|V_{n}\|\leq C\|\Delta(\theta_{\star})^{-1}\|\|\theta_{n}-\theta_{\star}\|.$ Define $A_{n}=C\Delta(\theta_{\star})^{-1}\|\theta_{n}-\theta_{\star}\|$, and conclude that $A_{n}$ is positive semi-definite with maximal eigenvalue $\lambda_{\text{max}}(A_{n})=C\|\theta_{n}-\theta_{\star}\|\lambda_{\text{min}}\{\Delta(\theta_{\star})\}\geq 0$. Therefore, over $\|t\|\leq\gamma_{n}$,
$-t^{\top}A_{n}t\leq t^{\top}V_{n}t\leq t^{\top}A_{n}t;$
applying the above into (7) yields
$-t^{\top}\Delta(\theta_{\star})t/2-t^{\top}A_{n}\Delta(\theta_{\star})t/2\leq Q_{n}(\theta_{n}+t/{\sqrt{n}})-Q_{n}(\theta_{n})$ $\leq-t^{\top}\Delta(\theta_{\star})t/2+t^{\top}A_{n}\Delta(\theta_{\star})t/2.$
Let $M^{\pm}_{n}=I\pm\Delta(\theta_{\star})^{-1/2}A_{n}\Delta(\theta_{\star})^{1/2}$, and rewrite the above as
$-\frac{1}{2}t^{\top}\Delta(\theta_{\star})^{1/2}M^{+}_{n}\Delta(\theta_{\star})^{1/2}t\leq Q_{n}(\theta_{n}+t/{\sqrt{n}})-Q_{n}(\theta_{n})$ $\leq-\frac{1}{2}t^{\top}\Delta(\theta_{\star})^{1/2}M^{-}_{n}\Delta(\theta_{\star})^{1/2}t.$
For $\|\theta_{n}-\theta_{\star}\|$ small enough, i.e., $n$ large enough, $\Delta(\theta_{\star})^{1/2}M^{\pm}_{n}\Delta(\theta_{\star})^{1/2}$ is positive-definite (with probability converging to one). Thus, for $n$ large we can bound the posterior probability over $\mathcal{N}_{\gamma_{n}}$ as
$\int_{\mathcal{N}_{\gamma_{n}}}\pi(\theta\mid\mathcal{Q}_{n})\mathrm{d}\theta$ $\leq\frac{\pi(\theta_{n}\mid\mathcal{Q}_{n})}{{n^{{d_{\theta}}/2}}}\left|\Delta(\theta_{\star})^{1/2}M_{n}^{-}\Delta(\theta_{\star})^{1/2}\right|^{-1/2}\int_{T^{-}_{n}}\exp(-x^{\top}x/2)\mathrm{d}x$
$\int_{\mathcal{N}_{\gamma_{n}}}\pi(\theta\mid\mathcal{Q}_{n})\mathrm{d}\theta$ $\geq\frac{\pi(\theta_{n}\mid\mathcal{Q}_{n})}{{n^{{d_{\theta}}/2}}}\left|\Delta(\theta_{\star})^{1/2}M_{n}^{+}\Delta(\theta_{\star})^{1/2}\right|^{-1/2}\int_{T^{+}_{n}}\exp(-x^{\top}x/2)\mathrm{d}x$
where
$T^{-}_{n}=\left\{x:\|x\|\leq\frac{\gamma_{n}{\sqrt{n}}}{\lambda_{\text{min}}\{\Delta(\theta_{\star})^{1/2}M_{n}^{-}\Delta(\theta_{\star})^{1/2}\}^{1/2}}\right\},\;T^{+}_{n}=\left\{x:\|x\|\leq\frac{\gamma_{n}{\sqrt{n}}}{\lambda_{\text{max}}\{\Delta(\theta_{\star})^{1/2}M_{n}^{+}\Delta(\theta_{\star})^{1/2}\}^{1/2}}\right\},$
and where we have used the fact that for any positive semi-definite matrix $M$ and $\gamma>0$
$\{x:\|x\|\leq\gamma/{\lambda_{\text{max}}(M)}^{1/2}\}\subseteq\{x:x^{\top}Mx\leq\gamma\}\subseteq\{x:\|x\|\leq\gamma/{\lambda_{\text{min}}(M)}^{1/2}\}.$
Under the restriction that $\gamma_{n}{\sqrt{n}}\rightarrow\infty$, $T_{n}^{+}$ and $T_{n}^{-}$ both converge to $\mathbb{R}^{d_{\theta}}$ and $\int_{T^{\pm}_{n}}\exp(-x^{\top}x/2)\mathrm{d}x\rightarrow{(2\pi)}^{d_{\theta}/2}$ as $n\rightarrow\infty$. Hence, with probability converging to one,
$|M_{n}^{+}|^{1/2}\int_{\mathcal{N}_{\gamma_{n}}}\pi_{n}(\theta\mid\mathcal{Q}_{n})\mathrm{d}\theta\leq\frac{\pi(\theta_{n}\mid\mathcal{Q}_{n})}{{n^{{d_{\theta}}/2}}}(2\pi)^{d_{\theta}/2}|\Delta^{-1}|^{1/2}$ $\leq|M_{n}^{-}|^{1/2}\int_{\mathcal{N}_{\gamma_{n}}}\pi(\theta\mid\mathcal{Q}_{n})\mathrm{d}\theta.$
Since $|M_{n}^{\pm}|\rightarrow 1$, $|\Delta|>0$ and $0\leq\int_{\mathcal{N}_{\gamma_{n}}}\pi(\theta\mid\mathcal{Q}_{n})\mathrm{d}\theta\leq 1$, with probability converging to one,
$0\leq\frac{\pi(\theta_{n}\mid\mathcal{Q}_{n})}{{n^{{d_{\theta}}/2}}}\leq 1/\{(2\pi)^{d_{\theta}}|\Delta^{-1}|\}^{1/2}.$
The following result is used in the proof of Lemma 1 and is a consequence of Proposition 1 in Chernozhukov and Hong, (2003).
Lemma 3.
Under the assumptions of Lemma 1, and for $R_{n}(\theta)$ as defined in the proof of Lemma 1, for each $\epsilon>0$ there exists a sufficiently small $\delta>0$ and $h>0$ large enough, such that
$\limsup_{n\rightarrow+\infty}P^{(n)}_{0}\left[\sup_{h/\sqrt{n}\leq\|\theta-\theta_{0}\|\leq\delta}\frac{|R_{n}(\theta)|}{1+{n}\|\theta-\theta_{0}\|^{2}}>\epsilon\right]<\epsilon$
and
$\limsup_{n\rightarrow+\infty}P^{(n)}_{0}\left[\sup_{\|\theta-\theta_{0}\|\leq h/\sqrt{n}}{|R_{n}(\theta)|}>\epsilon\right]=0.$
Proof.
The result is a specific case of Proposition 1 in Chernozhukov and Hong, (2003). Therefore, it is only necessary to verify that their sufficient conditions are satisfied in our context.
Assumptions (i)-(iii) in their result follow directly from Assumptions 1 and 2, and the normality of $\sqrt{n}\overline{m}_{n}(\theta_{\star})$ in Assumption 1. Their Assumption (iv) is stated as follows: for any $\epsilon>0$, there is a $\delta>0$ such that
$\limsup_{n\rightarrow+\infty}P^{(n)}_{0}\left\{\sup_{\|\theta-\theta^{\prime}\|\leq\delta}\frac{\sqrt{n}\|\{\overline{m}_{n}(\theta)-\overline{m}_{n}(\theta^{\prime})\}-\{\mathbb{E}\left[\overline{m}_{n}(\theta)\right]-\mathbb{E}\left[\overline{m}_{n}(\theta^{\prime})\right]\}\|}{1+n\|\theta-\theta^{\prime}\|}>\epsilon\right\}<\epsilon.$
In our context, this condition is satisfied by Assumption 1(iv). Hence, the result follows. ∎
A.4 Proofs of the Main Results
We first prove Theorem 1, as it simplifies the proof of Theorem 2.
Proof of Theorem 2.
The proof of Theorem 2 follows similar arguments to that of Theorem 1 in Frazier et al., (2021). Define $C_{\Delta}:=1/\{{(2\pi)}^{d_{\theta}}|\Delta(\theta_{\star})^{-1}|\}^{1/2}$ and $\Pi_{\star}(\gamma):=C_{\Delta}\int_{\|t\|\leq\gamma}\exp\left(-t^{\top}\Delta(\theta_{\star})t/2\right)\mathrm{d}t$. The result follows if $J_{n}=|\int_{\|t\|\leq\gamma}{\pi}(t\mid\mathcal{Q}_{n})\mathrm{d}t-\Pi_{\star}(\gamma)|=o_{p}(1).$
Rewrite the exact posterior $\pi(\theta\mid\mathcal{Q}_{n})$ as
$\pi(\theta\mid\mathcal{Q}_{n})$ $=\pi(\theta_{n}\mid\mathcal{Q}_{n})\frac{M_{n}(\theta)\pi(\theta)}{M_{n}(\theta_{n})\pi(\theta_{n})}\exp\left\{Q_{n}(\theta)-Q_{n}(\theta_{n})\right\},$ (8)
for any $\theta_{n}$ as in Lemma 1. For any $\gamma>0$, let $\gamma_{n}=\gamma/{\sqrt{n}}$, with $\gamma_{n}=o(1)$, and define $\mathcal{N}_{\gamma_{n}}=\{\theta\in\Theta:\|\theta-\theta_{n}\|\leq\gamma_{n}\}$. Plugging equation (6) in the proof of Lemma 2 into equation (8), we have
$\int_{\mathcal{N}_{\gamma_{n}}}\pi(\theta\mid\mathcal{Q}_{n})\mathrm{d}\theta=\pi(\theta_{n}\mid\mathcal{Q}_{n})\{1+o_{p}(1)\}\int_{\mathcal{N}_{\gamma_{n}}}\exp\left\{Q_{n}(\theta)-Q_{n}(\theta_{n})\right\}\mathrm{d}\theta.$ (9)
Now, consider the change of variables $\theta\mapsto t={\sqrt{n}}(\theta-\theta_{n})$, and the posterior $\pi(t\mid S_{n})=\pi(\theta_{n}+t/{\sqrt{n}}\mid\mathcal{Q}_{n})/{n^{{d_{\theta}}/2}}$. Noting that $\mathcal{N}_{\gamma_{n}}$ can be written as $\{t:\|t\|\leq\gamma\}$, we have
$\int_{\mathcal{N}_{\gamma_{n}}}\pi(\theta\mid\mathcal{Q}_{n})\mathrm{d}\theta$ $=\int_{\|t\|\leq\gamma}\pi(t\mid\mathcal{Q}_{n})\mathrm{d}t$
$=\frac{\pi(\theta_{n}\mid\mathcal{Q}_{n})}{{n^{{d_{\theta}}/2}}}\{1+o_{p}(1)\}\int_{\|t\|\leq\gamma}\exp\left\{Q_{n}(\theta_{n}+t/{\sqrt{n}})-Q_{n}(\theta_{n})\right\}\mathrm{d}t$
$\leq C_{\Delta}\{1+o_{p}(1)\}\int_{\|t\|\leq\gamma}\exp\left\{Q_{n}(\theta_{n}+t/{\sqrt{n}})-Q_{n}(\theta_{n})\right\}\mathrm{d}t;$ (10)
the second equality follows from equation (9) and the change of variables, and the last equation follows by Lemma 2 (with $P^{(n)}_{0}$-probability converging to one). Applying equation (10), we see that, up to an $o_{p}(1)$ term,
$J_{n}$ $\lesssim\left|\int_{\|t\|\leq\gamma}\left[\exp\left\{Q_{n}(\theta_{n}+{t}/{{\sqrt{n}}})-Q_{n}(\theta_{n})\right\}-\exp\left\{-t^{\top}\Delta(\theta_{\star})t/2\right\}\right]\mathrm{d}t\right|;$
the result follows if the right hand side of the above converges to zero in $P^{(n)}_{0}$-probability.
Using the same argument as in the proof of Lemma 2, for $A_{n}=C\Delta(\theta_{\star})^{-1}\|\theta_{n}-\theta_{\star}\|$, for some $C>0$, we have
$-t^{\top}\Delta(\theta_{\star})t/2-t^{\top}A_{n}\Delta(\theta_{\star})t/2\leq Q_{n}(\theta_{n}+t/{\sqrt{n}})-Q_{n}(\theta_{n})$ $\leq-t^{\top}\Delta(\theta_{\star})t/2+t^{\top}A_{n}\Delta(\theta_{\star})t/2.$
However, using the definition of $A_{n}$, over $\|t\|\leq\gamma$, the above equation simplifies to
$-t^{\top}\Delta(\theta_{\star})t/2-C\|\theta_{n}-\theta_{\star}\|\leq Q_{n}(\theta_{n}+t/{\sqrt{n}})-Q_{n}(\theta_{n})$ $\leq-t^{\top}\Delta(\theta_{\star})t/2+C\|\theta_{n}-\theta_{\star}\|.$ (11)
Using equations (10) and (11), and for $n$ large enough, we can bound the posterior over $\mathcal{N}_{\gamma}$ above and below as
$\frac{\pi(\theta_{n}\mid\mathcal{Q}_{n})}{{n^{{d_{\theta}}/2}}}\int_{\mathcal{N}_{\gamma}}\exp\left\{Q_{n}(\theta_{n}+t/{\sqrt{n}})-Q_{n}(\theta_{n})\right\}\mathrm{d}t$
$\leq\frac{\pi(\theta_{n}\mid\mathcal{Q}_{n})}{{n^{{d_{\theta}}/2}}}\exp(C\|\theta_{n}-\theta_{\star}\|)\int_{\mathcal{N}_{\gamma_{n}}}\exp\left\{-t^{\top}\Delta(\theta_{\star})t/2\right\}\mathrm{d}t$
$\geq\frac{\pi(\theta_{n}\mid\mathcal{Q}_{n})}{{n^{{d_{\theta}}/2}}}\exp(-C\|\theta_{n}-\theta_{\star}\|)\int_{\mathcal{N}_{\gamma_{n}}}\exp\left\{-t^{\top}\Delta(\theta_{\star})t/2\right\}\mathrm{d}t.$
As $n\rightarrow\infty$, it follows from Lemma 2, and the dominated convergence theorem that
$\Pi_{\star}(\gamma)\{1+o_{p}(1)\}$ $\leq\frac{\pi(\theta_{n}\mid\mathcal{Q}_{n})}{{n^{{d_{\theta}}/2}}}\int_{\mathcal{N}_{\gamma}}\exp\left\{Q_{n}(\theta_{n}+t/{\sqrt{n}})-Q_{n}(\theta_{n})\right\}\mathrm{d}t$
$\Pi_{\star}(\gamma)\{1+o_{p}(1)\}$ $\geq\frac{\pi(\theta_{n}\mid\mathcal{Q}_{n})}{{n^{{d_{\theta}}/2}}}\int_{\mathcal{N}_{\gamma}}\exp\left\{Q_{n}(\theta_{n}+t/{\sqrt{n}})-Q_{n}(\theta_{n})\right\}\mathrm{d}t.$
Hence, $J_{n}\rightarrow 0$ in $P^{(n)}_{0}$-probability.
Proof of Theorem 1.
Since $\theta_{\star}$ is assumed to be unique, we write $\Delta=\Delta(\theta_{\star})$. For some $\delta_{n}$ such that $\delta_{n}=o(1)$ and ${\sqrt{n}}\delta_{n}\rightarrow\infty$, split the region of integration into $\mathcal{T}_{n}=\mathcal{T}_{1n}\cup\mathcal{T}_{2n}$, where $\mathcal{T}_{1n}=\{h\leq\|t\|\leq\delta_{n}{\sqrt{n}}\}$ and $\mathcal{T}_{2n}=\{\|t\|\geq\delta_{n}{\sqrt{n}}\}$, for some arbitrary $h>0$. We consider the integral over each region separately.
Region $\mathcal{T}_{1n}$. Rewrite the posterior as
$\pi(\theta\mid\mathcal{Q}_{n})=\frac{\pi(\theta_{n}\mid\mathcal{Q}_{n})|M_{n}(\theta_{n})|^{1/2}}{\pi(\theta_{n})}|M_{n}(\theta)|^{-1/2}\exp\{Q_{n}(\theta)-Q_{n}(\theta_{n})\}\pi(\theta).$
Note that, from the definition of $\theta_{n}$ and Assumption 1(iv), $\|\sqrt{n}(\theta_{n}-\theta_{\star})\|=O_{p}(1)$ by Lemma 1. Under the uniqueness of $\theta_{\star}$, Lemma 2 implies
$\pi(\theta_{n}\mid\mathcal{Q}_{n})/{n^{{d_{\theta}}/2}}=C_{\Delta}+o_{p}(1)=C_{\Delta}=1/\{(2\pi)^{d_{\theta}}|\Delta^{-1}|\}^{1/2}+o_{p}(1).$ (12)
Over $\{\|t\|\leq\delta_{n}{\sqrt{n}}\}$, equation (12) and similar arguments to those used in the proof of Theorem 1 to obtain equation (10) yield
$\int_{\|\theta-\theta_{n}\|\leq\delta_{n}}\pi(\theta\mid\mathcal{Q}_{n})\mathrm{d}\theta$ $=\frac{\pi(\theta_{n}\mid\mathcal{Q}_{n})}{{n^{{d_{\theta}}/2}}}\{1+o_{p}(1)\}\int_{\|t\|\leq\delta_{n}{\sqrt{n}}}\pi(t\mid\mathcal{Q}_{n})\mathrm{d}t$
$=C_{\Delta}\{1+o_{p}(1)\}\int_{\mathcal{T}_{1n}}\exp\{Q_{n}(\theta_{n}+t/{\sqrt{n}})-Q_{n}(\theta_{n})\}\mathrm{d}t,$
and, up to an $o_{p}(1)$ term,
$\int_{\mathcal{T}_{1n}}\|t\||\pi(t\mid\mathcal{Q}_{n})-N\{t;0,\Delta^{-1}\}|\mathrm{d}t$
$=C_{\Delta}\int_{\mathcal{T}_{1n}}\|t\||\exp\{Q_{n}(\theta_{n}+{t}/{{\sqrt{n}}})-Q_{n}(\theta_{n})\}-\exp(-t^{\top}\Delta t/2)|\mathrm{d}t.$ (13)
Over $\mathcal{T}_{1n}$, arguments similar to those used in Theorem 1 to obtain equation (7) yield
$Q_{n}(\theta_{n}+t/{\sqrt{n}})-Q_{n}(\theta_{n})\leq-\frac{1}{2}t^{\top}\left(I-A_{n}\right)\Delta t+o_{p}(1),$ (14)
for $A_{n}$ as defined in that proof.
For some $0<h<\infty$, further split $\mathcal{T}_{1n}=\{\|t\|\leq h\}\cup\{h\leq\|t\|\leq\delta_{n}{\sqrt{n}}\}$, and consider first the set $\{\|t\|\leq h\}$. Now, recall that $A_{n}\rightarrow 0$ as $n\rightarrow\infty$, so that over $\|t\|\leq h$ equation (14) implies
$Q_{n}(\theta_{n}+t/{\sqrt{n}})-Q_{n}(\theta_{n})=-\frac{1}{2}t^{\top}\Delta t+o_{p}(1),$
and it follows that the integral in (13) is $o_{p}(1)$ over $\|t\|\leq h$.
Over $\{h\leq\|t\|\leq\delta_{n}{\sqrt{n}}\}$,
$\int_{h<\|t\|\leq{\sqrt{n}}\delta_{n}}\|t\|N\{t;0,\Delta^{-1}\}\mathrm{d}t$
can be made arbitrarily small by taking $h$ large enough and $\delta_{n}$ small enough, so that, applying (14), it suffices to show that for any $\varepsilon>0$ there exists an $h$ and $\delta_{n}$ such that, for some $n$ large enough,
$P^{(n)}_{0}\left[\int_{h<\|t\|\leq{\sqrt{n}}\delta_{n}}\|t\|\exp\{-t^{\top}\left(I-A_{n}\right)\Delta t/2\}\mathrm{d}t<\varepsilon\right]\geq 1-\varepsilon.$
However, for $h^{\prime}$ large enough, and all $h>h^{\prime}$, on the set $h<\|t\|\leq\delta_{n}{\sqrt{n}}$,
$\|t\|\exp(-t^{\top}\Delta t/2)=O(1/h).$
Therefore, for any $\varepsilon>0$, there is an $h$ large enough and a $\delta_{n}$ small enough such that
$\int_{h<\|t\|\leq{\sqrt{n}}\delta_{n}}\|t\|\exp(-t^{\top}\Delta t/2)\mathrm{d}t<\varepsilon.$
Since $A_{n}\rightarrow 0$, we can conclude that for some $n$ large enough, with $P^{(n)}_{0}$-probability at least $1-\varepsilon$,
$\int_{M<\|t\|\leq{\sqrt{n}}\delta}\|t\|\exp\{-t^{\top}\left(I-A_{n}\right)\Delta t/2\}\mathrm{d}t<\varepsilon.$
Region $\mathcal{T}_{2n}$. Again, $\int_{\mathcal{T}_{2n}}\|t\|N\{t;0,\Delta^{-1}\}\mathrm{d}t$ can be made arbitrarily small by taking $\delta_{n}{\sqrt{n}}$ large enough, and it remains to show that $\int_{\mathcal{T}_{2n}}\|t\|\pi(t\mid\mathcal{Q}_{n})=o_{p}(1)$.
Applying, in-turn, the expression for the exact posterior in (8), Assumptions 2, 3, and Lemma 2,
$\int_{\mathcal{T}_{2n}}\|t\|\pi(t\mid\mathcal{Q}_{n})\mathrm{d}t\leq$
$C\int_{\mathcal{T}_{2n}}{\|t\|M_{n}(\theta_{n}+t/n^{1/2})\pi(\theta_{n}+t/{\sqrt{n}})}\exp\{Q_{n}(\theta_{n}+t/{\sqrt{n}})-Q_{n}(\theta_{n})\}\mathrm{d}t,$
for some $C>0$ with probability converging to one. Using the change of variables $t\mapsto\theta$, the integral on the right hand side of the inequality is bounded by
$C\left\{1+o_{p}(1)\right\}{n^{(d_{\theta}+2)/2}}\int_{\|\theta-\theta_{\star}\|>\delta}\|\theta-\theta_{\star}\|\pi(\theta)M_{n}(\theta)\exp\{Q_{n}(\theta)-Q_{n}(\theta_{n})\}\mathrm{d}\theta,$ (15)
where the $o_{p}(1)$ term follows from the triangle inequality and consistency of $\theta_{n}$ for $\theta_{\star}$ (Lemma 1). For any $\delta>0$, and $Q(\theta)=-m(\theta)^{\top}W(\theta)^{-1}m(\theta)/2$, under Assumptions 1 and 2,
$\sup_{\|\theta-\theta_{\star}\|>\delta}n^{-1}\left\{Q_{n}(\theta)-Q_{n}(\theta_{n})\right\}$ $\leq\sup_{\|\theta-\theta_{\star}\|>\delta}\{Q(\theta)-Q(\theta_{\star})\}+o_{p}(1).$
From Assumption 2(iii), for any $\delta>0$ there exists some $\epsilon>0$ such that $\sup_{\|\theta-\theta_{\star}\|>\delta}\{Q(\theta)-Q(\theta_{\star})\}\leq-\epsilon.$ Therefore, for any $\delta>0$,
$\lim_{n\rightarrow\infty}P^{(n)}_{0}\left[\sup_{\|\theta-\theta_{\star}\|>\delta}\exp\{Q_{n}(\theta)-Q_{n}(\theta_{n})\}\leq\exp(-\epsilon n)\right]=1.$
Use the definition $M_{n}(\theta)=|W_{n}(\theta)^{-1}|^{1/2}$, and the above to obtain
$\int_{\mathcal{T}_{2n}}\|t\|\pi(t\mid\mathcal{Q}_{n})\mathrm{d}t$ $\leq C\{1+o_{p}(1)\}n^{(d_{\theta}+2)/2}\int_{\|\theta-\theta_{0}\|\geq\delta}M_{n}(\theta)\|\theta-\theta_{\star}\|\pi\left(\theta\right)\exp\{Q_{n}(\theta)-Q_{n}(\theta_{\star})\}\mathrm{d}\theta$
$\leq C\exp\left(-\epsilon n\right)n^{(d_{\theta}+2)/2}\int_{\|\theta-\theta_{\star}\|\geq\delta}M_{n}(\theta)\|\theta-\theta_{\star}\|\pi\left(\theta\right)\mathrm{d}\theta$
$\leq C\left\{\exp\left(-\epsilon n\right)n^{(d_{\theta}+2)/2}\right\}$
$=o_{p}(1),$
where the second inequality follows from the moment hypothesis in Assumption 3. ∎
Proof of Corollary 1.
For $\vartheta=\sqrt{n}(\theta-\theta_{n})$, consider the change of variables $\theta=\theta_{n}+\vartheta/\sqrt{n},$
$\overline{\theta}=\int_{\Theta}\theta\pi(\theta\mid\mathcal{Q}_{n})\mathrm{d}\theta=\int_{\mathcal{T}_{n}}(\theta_{n}+\vartheta/\sqrt{n})\pi(\vartheta\mid\mathcal{Q}_{n})\mathrm{d}\vartheta$
so that
$\sqrt{n}(\overline{\theta}-\theta_{n})=\int_{\mathcal{T}_{n}}\vartheta\pi(\theta\mid\mathcal{Q}_{n})\mathrm{d}\vartheta.$
When $\theta_{\star}$ is unique, standard show that $\theta_{n}:=\theta_{\star}-\Delta(\theta_{\star})^{-1}\mathcal{H}_{\star}^{\top}W(\theta_{\star})^{-1}\overline{m}_{n}(\theta_{\star})$. Hence, for $Z_{n}=\Delta^{-1}\mathcal{H}_{\star}^{\top}W(\theta_{\star})^{-1}\sqrt{n}\overline{m}_{n}(\theta_{\star})$, we have
$\sqrt{n}(\overline{\theta}-\theta_{n})=\sqrt{n}(\overline{\theta}-\theta_{\star})-Z_{n}=\int_{\mathcal{T}_{n}}\vartheta[\pi(\theta\mid\mathcal{Q}_{n})-N\{\vartheta;0,\Delta(\theta_{\star})^{-1}\}]\mathrm{d}\vartheta+\int_{\mathcal{T}_{n}}\vartheta N\{\vartheta;0,\Delta(\theta_{\star})^{-1}\}\mathrm{d}\vartheta.$
Since $\mathcal{T}_{n}\rightarrow\mathbb{R}^{d_{\theta}}$ as $n\rightarrow\infty$, $\int_{\mathcal{T}_{n}}\vartheta N\{\vartheta;0,\Delta(\theta_{\star})^{-1}\}\mathrm{d}\vartheta=o(1)$, and by Theorem 1,
$\left\|\int_{\mathcal{T}_{n}}\vartheta\pi(\vartheta\mid\mathcal{Q}_{n})\mathrm{d}\vartheta\right\|\leq\int_{\mathcal{T}_{n}}\|\vartheta\||\pi(\vartheta\mid\mathcal{Q}_{n})-N\{\vartheta;0,\Delta(\theta_{\star})^{-1}\}|\mathrm{d}\vartheta+o(1)=o_{p}(1).$
Therefore, we have that $\|\sqrt{n}(\overline{\theta}-\theta_{\star})-Z_{n}\|=o_{p}(1)$. By Assumption 1,
$Z_{n}\Rightarrow N\{0,\Delta(\theta_{\star})^{-1}\mathcal{H}_{\star}W(\theta_{\star})^{-1}\mathcal{I}_{\star}W(\theta_{\star})^{-1}\mathcal{H}_{\star}\Delta(\theta_{\star})^{-1}\}$
and the stated result follows. ∎
A.5 Conjugacy for Exponential Family Models in Natural Form
Suppose $Y_{1},\dots,Y_{n}\stackrel{{\scriptstyle iid}}{{\sim}}p_{\theta}(y)=\exp\{\eta(\theta)^{\top}S(y)-A(\theta)\}h(y),$ where $S:\mathcal{Y}\rightarrow\mathbb{R}^{d_{\theta}}$ is a vector of sufficient statistics, $h:\mathcal{Y}\rightarrow\mathbb{R}$ a reference measure or density on the sample space $\mathcal{Y}$, and $A:\Theta\rightarrow\mathbb{R}$ the log-partition function (see Lehmann and Casella, , 2006, Section 1.5 for further details). Then, the joint density $p_{\theta}^{(n)}(y)$ takes the form
$p_{\theta}^{(n)}(y)=\exp\left\{\eta(\theta)^{\top}\sum_{i=1}^{n}S(y_{i})-nA(\theta)\right\}\prod_{i=1}^{n}h(y_{i}),$
where $A(\theta)=\log\left[\int\exp\left\{\eta(\theta)^{\top}S(x)\right\}h(x)\mathrm{d}\mu(x)\right],$ and $\mu(x)$ is the Lebesgue measure.
Conducting inference on the natural parameter $\eta=\eta(\theta)$ is simplified by noting that if a conjugate prior is placed on $\eta$, then its Q-posterior has a closed-form expression. In particular, the above model has average scores $\overline{m}_{n}(\eta)=\nabla_{\eta}A(\eta)-Q_{n}$, where $S_{n}=n^{-1}\sum_{i=1}^{n}S(y_{i})$. Since $A(\eta)$ is non-random and the variance of $\overline{m}_{n}(\eta)$ can be estimated using the sample variance
$W_{n}:=\frac{1}{n}\sum_{i=1}^{n}\left\{S(y_{i})-S_{n}\right\}\left\{S(y_{i})-S_{n}\right\}^{\top},$
which does not depend on $\eta$. More generally, any consistent estimator of $\text{Cov}\{S(y_{i})\}$ can be used. One can then consider inference on $\eta$ using $\pi(\eta\mid\mathcal{Q}_{n})\propto\exp\{-nQ_{n}(\eta)\}\pi(\eta)$, where $Q_{n}(\eta)=\frac{1}{2}\left\{\nabla_{\eta}A(\eta)-S_{n}\right\}^{\top}W_{n}^{-1}\{\nabla_{\eta}A(\eta)-S_{n}\}.$
Define the mean parameter $\mu$ by the function $\mu=g(\eta)=\nabla_{\eta}A(\eta)$. In regular models the function $g(\eta)$ exists and is invertible for all $\eta$. The parameter $\mu=\mu(\eta)$ is referred to as the mean parameterization of the model and satisfies $\mu=\mathbb{E}_{Y\sim P^{(n)}_{\theta}}[Q_{n}]$. The form of the Q-posterior for $\eta$ then follows by finding the Q-posterior for $\mu$, and invoking a change of variables.
Lemma 4.
Suppose that $\mathcal{Y}=\mathbb{R}^{d}$. If the (transformed) prior beliefs for the mean parameter $\mu=g(\eta)$ is $\pi(\mu)\propto\exp\{-\frac{1}{2}(\mu-\mu_{0})^{\top}W_{0}^{-1}(\mu-\mu_{0})\}$, then the Q-posterior for $\eta$ is $\pi(\eta\mid\mathcal{Q}_{n})=N\{b_{n};g(\eta),\Sigma^{-1}_{n}\}|\nabla_{\eta}^{2}A(\eta)|$, where
$\Sigma^{-1}_{n}$ $=n^{-1}W_{0}\left[n^{-1}W_{n}+W_{0}\right]^{-1}W_{n},$
$b_{n}$ $=W_{0}\left[n^{-1}W_{n}+W_{0}\right]^{-1}S_{n}+W_{0}\left[n^{-1}W_{n}+W_{0}\right]^{-1}\frac{\mu_{0}}{n}.$
Lemma 4 demonstrates that the Q-posterior for the natural parameters $\eta$ is Gaussian if the prior for $\mu$ is Gaussian. Interestingly, calculating $|\nabla_{\eta}^{2}A(\eta)|$ can be avoided by first sampling $\widetilde{\mu}\sim N\{\mu;b_{n},\Sigma_{n}^{-1}\}$, and then (numerically) inverting the equation $\widetilde{\mu}=g(\eta)$ to obtain the draw $\widetilde{\eta}$. The latter is feasible when $g(\eta)=\nabla_{\eta}A(\eta)$ can be reliably calculated.
Proof of Lemma 4.
The Q-posterior under the mean parameterization $\mu=g(\eta)=\nabla_{\eta}A(\eta)$, with prior $\pi(\mu)\propto\exp\{-\frac{1}{2}(\mu-\mu_{0})^{\top}W_{0}^{-1}(\mu-\mu_{0})\}$, where $\mu_{0}$ and $W_{0}$ are known hyper-parameters. Define $\bar{S}_{n}=n^{-1}Q_{n}$, and note
$n^{-1}m_{n}(\eta)=g(\eta)-\bar{S}_{n}=\mu-\bar{S}_{n}=n^{-1}m_{n}(\mu).$
Hence, writing $\bar{S}_{n}=n^{-1}Q_{n}$, the Q-posterior for $\mu$ is
$\pi(\mu\mid\mathcal{Q}_{n})\propto\exp\left\{-\frac{n}{2}\left(\mu-\bar{S}_{n}\right)^{\top}W_{n}^{-1}\left(\mu-\bar{S}_{n}\right)\right\}\exp\left\{-\frac{1}{2}(\mu-\mu_{0})^{\top}W_{0}^{-1}(\mu-\mu_{0})\right\}.$
We now show that $\pi(\mu\mid\mathcal{Q}_{n})=N\{\mu;b_{n},\Sigma_{n}^{-1}\}$. With some algebra,
$\exp\left\{-\frac{1}{2}\mu^{\top}\left[(W_{n}/n)^{-1}+W_{0}^{-1}\right]\mu-\mu^{\top}\left[(W_{n}/n)^{-1}\bar{S}_{n}+\mu_{0}\right]\right\}$
$=\exp\left\{-\frac{1}{2}\mu^{\top}\left[(W_{n}/n)^{-1}+W_{0}^{-1}\right]\mu-\mu^{\top}\left[(W_{n}/n)^{-1}+W_{0}^{-1}\right]\left[(W_{n}/n)^{-1}+W_{0}^{-1}\right]^{-1}\left[(W_{n}/n)^{-1}\bar{S}_{n}+\mu_{0}\right]\right\}$
$\propto\exp\bigg{\{}-\frac{1}{2}\left(\mu-\Sigma_{n}^{-1}\left[(W_{n}/n)^{-1}\bar{S}_{n}+\mu_{0}\right]\right)^{\top}\Sigma_{n}\left(\mu-\Sigma_{n}^{-1}\left[(W_{n}/n)^{-1}\bar{S}_{n}+\mu_{0}\right]\right)\bigg{\}}$
$=\exp\left\{-\frac{1}{2}(\mu-b_{n})^{\top}\Sigma_{n}(\mu-b_{n})\right\}$
so that
$\Sigma_{n}^{-1}$ $=\left[(W_{n}/n)^{-1}+W_{0}^{-1}\right]^{-1}=W_{0}\left[n^{-1}W_{n}+W_{0}\right]^{-1}W_{n}\frac{1}{n},$
where the second equality follows from the Woodbury identity, and
$b_{n}$ $=\Sigma_{n}^{-1}\left[(W_{n}/n)^{-1}\bar{S}_{n}+\mu_{0}\right]$
$=W_{0}\left[n^{-1}W_{n}+W_{0}\right]^{-1}\bar{S}_{n}+W_{0}\left[n^{-1}W_{n}+W_{0}\right]^{-1}W_{n}\frac{\mu_{0}}{n}\mu_{0}.$
Hence, $\pi(\mu\mid\mathcal{Q}_{n})=N\{\mu;b_{n},\Sigma_{n}^{-1}\}$. For a regular exponential family, the parameter change from $\mu\mapsto\eta=g^{-1}(\mu)$ exists if the model is identifiable (in $\eta$). A change of variables $\mu\mapsto\eta$ then implies
$\pi(\eta\mid\mathcal{Q}_{n})=\pi\{g(\eta)\mid\mathcal{Q}_{n}\}|\nabla_{\eta}g(\eta)|=N\{g(\eta);b_{n},\Sigma_{n}^{-1}\}|\nabla_{\eta}^{2}A(\eta)|,$
where the second equality follows since $g(\eta)=\nabla_{\eta}A(\eta)$. ∎